### v1

# An abstract framework for heterogeneous coupling: stability, approximation and applications

Abstract.
Introducing a coupling framework reminiscent of FETI methods, but here on abstract form, we establish conditions for stability and minimal requirements for well-posedness on the continuous level, as well as conditions on local solvers for the approximation of subproblems. We then discuss stability of the resulting Lagrange multiplier methods and show stability under a mesh conditions between the local discretizations and the mortar space. If this condition is not satisfied we show how a stabilization, acting only on the multiplier can be used to achieve stability. The design of preconditioners of the Schur complement system is discussed in the unstabilized case. Finally we discuss some applications that enter the framework.
Key words and phrases:
Coupling methods; FETI; domain decomposition; FEM/BEM coupling; mixed dimensional problems
1991 Mathematics Subject Classification:
65N55
1. Introduction
The efficient approximation of linear problems set in infinite dimensional spaces is becoming increasingly important. A key example is the numerical approximation of solutions to partial differential equations (PDE) or integral equations, that are omni present in science and technology. Discretization results in linear systems that become increasingly large as the scale and the complexity of the problem grow. In view of the increasing availability of high performance computing environments, it is then advantageous to resort to solution strategies leveraging, possibly iteratively, local solvers for sets of (suitably coupled) subproblems.
The common approach in this framework is to discretize the global continuous problem into a possibly huge algebraic system, and then split the latter as a system of coupled algebraic subsystems. In this case the coupling condition often consists in the identification of unknowns in the subsystems corresponding to the same global unknown. An alternative approach, which is lately gaining increasing interest, is, instead, to decompose the original problem already at the continuous/infinite dimensional level, thus obtaining a system of coupled infinite dimensional subproblems, to be successively discretized. This is the approach that underlies a number of non conforming domain decomposition methods, such as the mortar method [4, 1, 31], the three fields domain decomposition method [11, 7], or the domain decomposition method based on local Dirichlet–Neumann maps proposed in [25], which all start from a domain decomposition of the continuous problem.
A key feature of this approach is that the subproblems can use different models and be discretized independently of each other using bespoke solvers that communicate only through the coupling conditions. This offers the possibility of resorting to independent implementations for the solvers of the different subproblems, an appealing feature, particularly in the framework of multiphysics and multiscale simulations, where highly specialized and optimized subproblem solvers might be available, and where the whole coupled problem might be excessively complex. Indeed, in this framework, we see a growing interest around the different issues related to “code coupling” [22, 21, 24], and the development of different computational environments aimed at enabling the interaction of existing solvers, individually simulating different portions of a larger complex physical problems (see, for instance, [28, 12, 30]).
The aim of this paper is to provide a theoretical foundation for such approaches where heterogeneous large scale problems are solved using different highly optimized codes to handle the different subproblems. Therefore we study, in an abstract framework, with problem, subproblems and coupling that are not a priori associated to a physical space and model, a strategy which is agnostic to the different solvers, assuming only that they produce a sufficiently accurate approximation of their target quantities. Following an approach inspired by the widely used Finite Element Tearing and Interconnecting (FETI) method [20, 18, 19, 23], by formally solving the local problems, up to a possible non-trivial, but finite dimensional, kernel of the local operator, the system is reduced to a problem where only the unknowns playing the role of Lagrangian multipliers, enforcing the coupling between subproblems, are to be computed. While the FETI method was developed in the framework of conforming domain decomposition for PDEs, we believe a large class of other problems, including multiscale models [17], FEM/BEM coupling [6], or PDEs set on lower dimensional manifolds connected in a network, can benefit from it and from the related techniques developed to handle the resulting algebraic system.
Our abstract starting point allows us to remain agnostic as to the local approximation strategy to the furthest possible extent: discretization of the local subproblem consists, in our abstract famework, in selecting a non-specified “local solver”. This important feature of our approach allows the coupling of highly optimized codes with solvers that use different methods, for instance finite difference methods, finite element methods, spectral methods or boundary element methods. By carrying out our analysis in an abstract framework, independent of the discretization methodology, we aim to establish the weakest conditions that such solvers must satisfy in order to guarantee the well posedness of the resulting global method. To establish such minimal conditions we analyze the properties of the reduced coupling system, for which well-posedness is proven at the continuous level. This continuous framework is then exploited when local discretizations are introduced, which we treat as “black boxes”, and conditions are proposed that are sufficient (but not necessary) to establish uniform stability of the discretization. To cater for the situation where such conditions, which implicitely enforce a compatibility condition between local solvers and multiplier discretization, can not be satisfied or lead to inconvenient discretizations, we introduce stabilization methods in the spirit of [10], generalizing the ideas of [15, 14]. We establish optimal error estimates for the discretization and we discuss the design of preconditioners for the unstabilized reduced system. Finally we discuss some different examples that enter the framework.
2. Problem setting
We consider an abstract problem of the form
[…]
where $\mathscr{V}$ is a separable Hilbert space, $\mathscr{A}:\mathscr{V}\to\mathscr{V}^{\prime}$ denotes a boundedly invertible linear operator, and where $f\in\mathscr{V}^{\prime}$ is the given right hand side. We assume that such a problem can be split as a system of “local” subproblems on smaller Hilbert spaces $V_{k}$, $k=1,\ldots,N$ subject to a coupling constraint. More precisely, we assume that we are given injective “restriction” operators $\mathscr{R}_{k}:\mathscr{V}\to V_{k}$, $k=1,\cdots,N$, and we define the composite operator $\mathscr{R}$ which maps $\mathscr{V}$ into $\prod_{k=1}^{N}V_{k}$, as
$\mathscr{R}=\left(\begin{array}[]{c}\mathscr{R}_{1}\\ \vdots\\ \mathscr{R}_{N}\end{array}\right).$
We then assume that there exist operators $B_{k}:V_{k}\to\Lambda^{\prime}$, $\Lambda$ being a second separable Hilbert space, such that the operator $\mathscr{R}$ is an isomorphism of $\mathscr{V}$ onto the space $\widehat{V}\subseteq V=\prod_{k=1}^{N}V_{k}$, defined as
$\widehat{V}=\{v=(v_{k})_{k}\in V:\sum_{k}B_{k}v_{k}=0\ \text{ in }\Lambda^{\prime}\}.$
We let $b_{k}:V_{k}\times\Lambda\to\mathbb{R}$ denote the bilinear form corresponding to $B_{k}$:
$b_{k}(u,\mu)=\langle B_{k}u,\mu\rangle,\qquad\forall\mu\in\Lambda.$
[…]
Assume now that continuous bilinear forms $a_{k}:V_{k}\times V_{k}\to\mathbb{R}$ and linear operators $f_{k}\in V_{k}^{\prime}$ are given, such that for all $\mathsf{v},\mathsf{w}\in\mathscr{V}$ it holds that
$\langle\mathscr{A}\mathsf{v},\mathsf{w}\rangle=\sum_{k=1}^{N}a_{k}(\mathscr{R}_{k}\mathsf{v},\mathscr{R}_{k}\mathsf{w}),\qquad\langle\mathsf{f},\mathsf{w}\rangle=\sum_{k=1}^{N}\langle f_{k},\mathscr{R}_{k}\mathsf{w}\rangle,$
and let $A_{k}:V_{k}\to V_{k}^{\prime}$ denote the linear operator corresponding to $a_{k}$:
$\langle A_{k}u,v\rangle=a_{k}(u,v),\qquad\forall u,v\in V_{k}.$
We assume that $\ker A_{k}$ is finite dimensional and coincides with $\ker A_{k}^{T}$, and that $a_{k}$ is coercive on the subspace
$C_{k}=(\ker A_{k})^{\perp}\subset V_{k}.$
Letting
$A=\left[\begin{array}[]{ccc}A_{1}&0&0\\ 0&\ddots&0\\ 0&0&A_{N}\end{array}\right],\qquad B=[B_{1},\cdots,B_{N}],\qquad f=\left[\begin{array}[]{c}f_{1}\\ \vdots\\ f_{N}\end{array}\right],$
we can finally rewrite Problem 2.1 in the following saddle point form
Problem 2.1.
Find $u\in V$ and $\lambda\in\Lambda$ such that
(2.2) $Au-B^{T}\lambda$ $=$ $f,$
(2.3) $Bu$ $=$ $0$
Letting $\|\cdot\|_{V}$ denote the product norm in $V$, we assume that an inf-sup condition for $B$ holds, of the form [^1] Here and in the following we use the notation $x\lesssim y$ to indicate that the quantity $x$ is lower that the quantity $y$ times a positive constant independent on any relevant parameter, particularly the ones related to the forthcoming discretizations.
(2.4) $\inf_{\lambda\in\Lambda}\sup_{v=(v^{k})\in V}\frac{b(v,\lambda)}{\|v\|_{V}\|\lambda\|_{\Lambda}}\gtrsim 1,\qquad\text{ where }b(v,\lambda)=\sum_{k}b_{k}(v^{k},\lambda).$
We do not assume that $A_{k}$ is invertible for all $k$, however the well posedness of equation (2.1), together with the above inf-sup condition, implies the well posedness of Problem 2.1.
In order to define the numerical method aimed at solving our problem, we follow, already at the continuous level, the approach underlying the FETI domain decomposition method [20]. We start by eliminating the unknown $u$. To this aim for all $k$ we introduce a bounded linear pseudo inverse $A^{+}_{k}$ of $A_{k}$, $A_{k}^{+}:V_{k}^{\prime}\to C_{k}\subset V_{k}$, defined as follows
(2.5) $a_{k}(A^{+}_{k}g,w)=\langle g,w\rangle,\qquad\text{ for all }w\in C_{k}.$
Letting $\mathfrak{R}(A)$ denote the range of $A$, such a pseudo inverse is well defined and verifies
$A_{k}A_{k}^{+}g=g,\quad\text{ for all }g\in\mathfrak{R}(A_{k}),\qquad\text{and}\qquad\|A_{k}^{+}Av\|_{V_{k}}\lesssim\|A_{k}v\|_{V^{\prime}_{k}},\qquad\forall v\in V_{k}.$
We then assemble the linear operator $A^{+}:V^{\prime}\to V$ as
$A^{+}=\left[\begin{array}[]{ccc}A^{+}_{1}&0&0\\ 0&\ddots&0\\ 0&0&A^{+}_{N}\end{array}\right].$
Letting the bilinear operator $a:V\times V\to\mathbb{R}$ be defined as
$\qquad a(u,v)=\langle Au,v\rangle=\sum_{k}a_{k}(u^{k},v^{k}),$
we can easily see that we have
$a(A^{+}Av,A^{+}Av)\gtrsim\|A^{+}Av\|_{V}^{2},\qquad\text{ for all }v\in V.$
Equation (2.2) trivially implies that $f+B^{T}\lambda=Au$. Then, letting
$Z=\ker A\subset V,$
we immediately see that for some unknown $z\in Z$, the solution $(u,\lambda)$ of Problem 2.1 satisfies
(2.6) $u=A^{+}(f+B^{T}\lambda)+z.$
Plugging this expression in (2.3) we obtain an identity in $\Lambda^{\prime}$ of the form
$BA^{+}B^{T}\lambda+Bz=-BA^{+}f.$
If we test (2.2) with any $z\in Z$, observing that, since, by the assumption that $\ker A=\ker A^{T}$, we have
$\langle Au,z\rangle=\langle u,A^{T}z\rangle=0,$
we immediately obtain that
(2.7) $\langle B^{T}\lambda+f,z\rangle=0,\ \forall z\in Z.$
Letting $G:Z\to\Lambda^{\prime}$ denote the restriction to $Z$ of $B$, we can rewrite (2.7) as an identity in $Z^{\prime}$:
$G^{T}\lambda+f=0,$
where, by abuse of notation, we let $f\in Z^{\prime}$ denote the restriction of $f$ to $Z$.
In other words we transformed the original saddle point problem to the following reduced saddle point problem, whose unknowns are the multiplier $\lambda$ and the $\ker A$ component $z$ of the solution.
Problem 2.2.
Given $f\in V^{\prime}$ find $\lambda\in\Lambda$ and $z\in Z$ such that
$S\lambda+Gz$ $=$ $g,$
$G^{T}\lambda$ $=$ $-f,$
where $S=BA^{+}B^{T}$ and $g=-BA^{+}f$.
It will be convenient, in the following, to write Problem 2.2 in weak form as: find $\lambda\in\Lambda$ and $z\in Z$ such that
$s(\lambda,\mu)+b(z,\mu)$ $=$ $\langle g,\mu\rangle,\qquad\forall\mu_{\delta}\in\Lambda_{\delta},$
$b(w,\lambda)$ $=$ $-\langle f,w\rangle\qquad\forall w\in Z.$
with $s:\Lambda\times\Lambda\to\mathbb{R}$ defined by
$s(\lambda,\mu)=\langle S\lambda,\mu\rangle.$
Note that the presence of a nontrivial kernel $Z$ indicates that a global coupling between the subproblems remains. The cardinality of the finite dimensional set $Z$ is in this sense a measure of how efficient the decoupling in local subproblems is. We will assume below that the dimension of $Z$ is relatively small, implying that the decomposition at hand is efficient.
This new saddle point problem, defined already at the continuous level, is the starting point of our abstract coupling algorithm. We start by verifying the well posedness of such a problem. We have the following Theorem.
Theorem 2.3.
The following inf-sup condition holds
(2.8) $\inf_{z\in Z}\sup_{\lambda\in\Lambda}\frac{\langle Bz,\lambda\rangle}{\|z\|_{V}\|\lambda\|_{\Lambda}}=\inf_{z\in Z}\sup_{\lambda\in\Lambda}\frac{b(z,\lambda)}{\|z\|_{V}\|\lambda\|_{\Lambda}}\gtrsim 1.$
Moreover the operator $s$ satisfies
(2.9) $s(\lambda,\lambda)\gtrsim\|\lambda\|^{2}_{\Lambda}\quad\text{ for all }\lambda\in\widehat{\Lambda}=\ker G^{T}=\{\lambda\in\Lambda:\ b(z,\lambda)=0\ \forall z\in Z\}$
and
(2.10) $s(\lambda,\mu)\lesssim\|\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
Proof.
We prove (2.8) by contradiction: more precisely we assume that for each $\epsilon>0$ there exist $z_{\epsilon}\in Z$ with $\|z_{\epsilon}\|_{V}=1$ such that for all $\lambda$ with $\|\lambda\|_{\Lambda}=1$ it holds that $b(z_{\epsilon},\lambda)<\epsilon$. We can then extract a weakly convergent sequence $z_{n}$. Let $z$ denote the weak limit of the sequence: $z_{n}\rightharpoonup z$. By definition of weak convergence we immediately see that $z\in\ker B$. Now we recall that $\ker B\cap Z=\ker B\cap\ker A=\{0\}$. As $Z$ is finite dimensional, this is in contradiction with $\|z_{n}\|_{V}=1$.
We next prove the coercivity of $s$ on $\widehat{\Lambda}$. We start by observing that $B^{T}\widehat{\Lambda}\subseteq\mathfrak{R}(A)$. Indeed, as, by assumption, $a$ is coercive in $\mathfrak{R}(A^{+})$, there exist $w\in\mathfrak{R}(A^{+})$ such that for all $v\in\mathfrak{R}(A^{+})$ it holds that
$a(w,v)=\langle B^{T}\lambda,v\rangle,\qquad\|w\|_{V}\lesssim\|B^{T}\lambda\|_{V^{\prime}}\lesssim\|\lambda\|_{\Lambda}.$
We claim that $Aw=B^{T}\lambda$ in $V^{\prime}$. Indeed, for all $v\in V$ we have $v=A^{+}Av+z$ with $z\in Z$, so that we can write
(2.11) $\langle Aw,v\rangle=a(w,A^{+}Av+z)=a(w,A^{+}Av)=\langle B^{T}\lambda,A^{+}Av\rangle+\langle G^{T}\lambda,z\rangle=\langle B^{T}\lambda,v\rangle.$
Now, given $\lambda\in\widehat{\Lambda}$ we have
$s(\lambda,\lambda)=\langle BA^{+}B^{T}\lambda,\lambda\rangle=\langle A^{+}B^{T}\lambda,B^{T}\lambda\rangle.$
On the other hand, as $B^{T}\lambda\in\mathfrak{R}(A)$, we have
$\|B^{T}\lambda\|_{V^{\prime}}=\|AA^{+}B^{T}\lambda\|_{V^{\prime}}\lesssim\|A^{+}B^{T}\lambda\|_{V}.$
Then we can write
$s(\lambda,\lambda)=a(A^{+}B^{T}\lambda,A^{+}B^{T}\lambda)\gtrsim\|A^{+}B^{T}\lambda\|_{V}^{2}\gtrsim\|B^{T}\lambda\|_{V^{\prime}}^{2}.$
We have, thanks to the inf-sup condition (2.4)
$\|B^{T}\lambda\|_{V^{\prime}}=\sup_{v\in V}\frac{b(v,\lambda)}{\|v\|_{V}}\gtrsim\|\lambda\|_{\Lambda},$
which allows us to conclude that
$s(\lambda,\lambda)\gtrsim\|\lambda\|^{2}_{\Lambda}.$
The continuity (2.10) follows, since
$s(\lambda,\mu)=\langle A^{+}B^{T}\lambda,B^{T}\mu\rangle\lesssim\|A^{+}B^{T}\lambda\|_{V}\|A^{+}B^{T}\mu\|_{V}\lesssim\|\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
Classical results on saddle point problems [9], together with the well posedness of Problem 2.1, imply the following corollary.
Corollary 2.4.
Problem 2.2 admits a unique solution $(\lambda,z)$ such that, letting $\mathsf{u}$ denote the solution of (2.1), we have
$\mathscr{R}\mathsf{u}=A^{+}(B^{T}\lambda+f)+z.$
3. Discretization
We discretize Problem 2.2 by a non conforming Galerkin discretization, involving both approximating the multiplier space $\Lambda$ with a finite dimensional space $\Lambda_{\delta}$, and numerically evaluating the pseudo inverses $A_{k}^{+}$. More precisely, for all $k$ we consider approximate linear solvers $A^{+}_{h,k}:V^{\prime}\to C_{k}$, that we will treat as a black-boxes. We then let
$A^{+}_{h}=\left[\begin{array}[]{ccc}A^{+}_{h,1}&0&0\\ 0&\ddots&0\\ 0&0&A^{+}_{h,N}\end{array}\right],$
and we define an approximate bilinear form $s_{h}:\Lambda\times\Lambda\to\mathbb{R}$ and an approximate right hand side $g_{h}$ as
$s_{h}(\lambda,\mu)=\langle BA^{+}_{h}B^{T}\lambda,\mu\rangle,\qquad g_{h}=-BA^{+}_{h}f.$
Let $\Lambda_{\delta}\subseteq\Lambda$ be a finite dimensional subspace depending on a parameter $\delta>0$, playing the role of a meshsize. Since the space $Z=\ker A$ is, by assumption, finite dimensional, we do not need to discretize it. We then consider the following discrete problem.
Problem 3.1.
Find $\lambda_{\delta}\in\Lambda_{\delta}$ and $z^{*}\in Z$ such that
$s_{h}(\lambda_{\delta},\mu_{\delta})+b(z^{*},\mu_{\delta})$ $=$ $\langle g_{h},\mu_{\delta}\rangle,\qquad\forall\mu_{\delta}\in\Lambda_{\delta},$
$b(w,\lambda_{\delta})$ $=$ $-\langle f,w\rangle,\qquad\forall w\in Z.$
As we want to treat them as black boxes, we want the assumptions on the approximate pseudo inverses to be as weak as possible. We want to understand under which minimal conditions Problem 3.1 is an approximation to Problem 2.2. To this aim it is only natural to require that the black box approximate linear solvers $A^{+}_{h,k}$ are approximations of the continuous pseudo inverses $A^{+}_{k}$, at least when the data have some extra smoothness.
To quantify, in our abstract setting, the concept of “extra smoothness”, we introduce infinite dimensional subspaces of $V_{k}$ and $\Lambda$, playing the role of higher regularity spaces. More precisely we let $W^{r}_{k}\subseteq V_{k}$ and $\Xi^{r}\subseteq\Lambda$, $r\geq 0$, denote a scale of subspaces of $V_{k}$ and $\Lambda$ respectively, such that $W^{0}_{k}=V_{k}$ and $\Xi^{0}=\Lambda$, and such that, for $t<r$, $\Xi^{r}\subset\Xi^{t}$ and $W^{r}_{k}\subset W_{k}^{t}$. We assume that, in this smoothness scale, $A^{+}B^{T}$ is smoothness preserving up to a certain range $r_{0}$, that is, that $\lambda\in\Xi^{r}$ implies that $A^{+}B^{T}\lambda\in W^{r}=\prod_{k}W^{r}_{k}$, and
(3.1) $\|A^{+}B^{T}\lambda\|_{W^{r}}\leq C_{r}\|\lambda\|_{\Xi^{r}},\qquad 0\leq r\leq r_{0}.$
We make one of the two following assumptions on the behavior of the black-box solver $A_{h}^{+}$ as the “meshsize” parameter $h$ tends to zero.
Assumption 3.2.
The discrete solver $A^{+}_{h}$ satisfies the following estimate: $A^{+}f\in W^{r}$, $0\leq r\leq R$, implies
$\|(A^{+}_{h}-A^{+})f\|_{V}\lesssim h^{r}\|A^{+}f\|_{W^{r}}.$
Assumption 3.3.
The discrete solver $A^{+}_{h}$ satisfies the following estimate: $A^{+}f\in W^{r}$, $0\leq r\leq R$, implies
$\|B(A^{+}_{h}-A^{+})f\|_{\Lambda^{\prime}}\lesssim h^{r}\|A^{+}f\|_{W^{r}}.$
Assumption 3.2 ensures the individual convergence of approximations for all the subproblems, whereas the weaker Assumption 3.3 only ascertains that the local solvers are sufficiently accurate to approximate the coupling. If only the latter holds, some other local solvers must be eventually used to reconstruct the full solution of the subproblems from the computed $\lambda_{\delta}$ and $z^{*}$. Note also that Assumption 3.2 with $r=0$ reduces to the boundedness of $A^{+}_{h}$.
We next make the quite natural assumption that $\Lambda_{\delta}$ is an order $m$ approximation space.
Assumption 3.4.
For all $\lambda\in\Xi^{r}$, $r>0$,
$\inf_{\mu_{\delta}\in\Lambda_{\delta}}\|\lambda-\mu_{\delta}\|_{\Lambda}\lesssim\delta^{\min\{r,m\}}\|\lambda\|_{\Xi^{r}}.$
The analysis below also requires the following inverse inequality to hold true in $\Lambda_{\delta}$.
Assumption 3.5.
For some $t>0$, $\Lambda_{\delta}\subset\Xi^{t}$, and for all $\lambda_{\delta}\in\Lambda_{\delta}$,
(3.2) $\|\lambda_{\delta}\|_{\Xi^{r}}\lesssim\delta^{-r}\|\lambda_{\delta}\|_{\Lambda},\quad 0<r\leq t.$
We have the following lemma.
Lemma 3.6.
Under Assumptions 3.3 and 3.5, the bilinear form $s_{h}$ is continuous on $\Lambda\times\Lambda_{\delta}$, and there exists a constant $C_{\Lambda}$ such that if $h/\delta<C_{\Lambda}$ then $s_{h}$ is coercive on $\widehat{\Lambda}_{\delta}=\Lambda_{\delta}\cap\widehat{\Lambda}$.
Proof.
We start by observing that, for $r>0$ sufficiently small, for $\lambda\in\Lambda$ and $\mu_{\delta}\in\Lambda_{\delta}$, we can write
(3.3) $\langle B(A^{+}-A^{+}_{h})B^{T}\lambda,\mu_{\delta}\rangle\lesssim\|B(A^{+}-A^{+}_{h})B^{T}\lambda\|_{\Lambda^{\prime}}\|\mu_{\delta}\|_{\Lambda}\\ \lesssim h^{r}\|A^{+}B^{T}\lambda\|_{W^{r}}\|\mu_{\delta}\|_{\Lambda}\lesssim h^{r}\|\lambda\|_{\Xi^{r}}\|\mu_{\delta}\|_{\Lambda}.$
By taking $r=0$ in (3.3) it follows that
$s_{h}(\lambda,\mu_{\delta})=s(\lambda,\mu_{\delta})+\langle B(A^{+}_{h}-A^{+})B^{T}\lambda,\mu_{\delta}\rangle\lesssim\|\lambda\|_{\Lambda}\|\mu_{\delta}\|_{\Lambda},$
which proves the continuity. Now, note that if $\lambda=\lambda_{\delta}\in\Lambda_{\delta}$ in (3.3), then by Assumption 3.5 we have, for some $r>0$,
$h^{r}\|\lambda_{\delta}\|_{\Xi^{r}}\|\mu_{\delta}\|_{\Lambda}\lesssim h^{r}\delta^{-r}\|\lambda_{\delta}\|_{\Lambda}\|\mu_{\delta}\|_{\Lambda},$
that is, for some constant $C>0$ we have
$\langle B(A^{+}-A^{+}_{h})B^{T}\lambda_{\delta},\lambda_{\delta}\rangle\leq Ch^{r}\delta^{-r}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
Let now $\lambda_{\delta}\in\widehat{\Lambda}_{\delta}=\Lambda_{\delta}\cap\widehat{\Lambda}$. As $\lambda_{\delta}\in\widehat{\Lambda}$, we know, by Theorem 2.3, that $s(\lambda_{\delta},\lambda_{\delta})\geq c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}$ for some $c_{s}>0$. We can write
$s_{h}(\lambda_{\delta},\lambda_{\delta})=\langle BA^{+}_{h}B^{T}\lambda_{\delta},\lambda_{\delta}\rangle=\langle BA^{+}B^{T}\lambda_{\delta},\lambda_{\delta}\rangle-\langle B(A^{+}-A^{+}_{h})B^{T}\lambda_{\delta},\lambda_{\delta}\rangle=$
$s(\lambda_{\delta},\lambda_{\delta})-\langle B(A^{+}-A^{+}_{h})B^{T}\lambda_{\delta},\lambda_{\delta}\rangle\geq c_{s}\|\lambda_{\delta}\|^{2}_{\Lambda}-\langle B(A^{+}-A^{+}_{h})B^{T}\lambda_{\delta},\lambda_{\delta}\rangle.$
We conclude that
$s_{h}(\lambda_{\delta},\lambda_{\delta})\geq c_{s}\|\lambda_{\delta}\|^{2}_{\Lambda}-Ch^{r}\delta^{-r}\|\lambda_{\delta}\|_{\Lambda}^{2},$
which, provided $h$ and $\delta$ are chosen in such a way that $h^{r}\delta^{-r}<c_{s}/C$, implies that $s_{h}$ is coercive. ∎
Finally, for the well posedness of the discrete problem we will need an inf-sup condition allowing to control $Z$ with elements of $\Lambda_{\delta}$. As $Z$ is a small fixed space (independent of $\delta$) the following mild assumption will be sufficient to this aim.
Assumption 3.7.
There is a (small) finite dimensional subspace $\Lambda_{0}\subset\Lambda$ such that $\Lambda_{0}\subseteq\cap_{\delta}\Lambda_{\delta}$ and such that
$\inf_{z\in Z}\sup_{\mu\in\Lambda_{0}}\frac{b(z,\mu)}{\|z\|_{V}\|\mu\|_{\lambda}}\gtrsim 1.$
Assumption 3.7 combined with Lemma 3.6 ensure the well posedness of Problem 3.4, which yields an optimal approximation to the solution of the original problem, as stated by the following theorem.
Theorem 3.8.
Under Assumptions 3.3, 3.5 and 3.7, $u-z\in W^{r}$, $r<R$ implies
$\|\lambda-\lambda_{\delta}\|_{\Lambda}+\|z-z^{*}\|_{V}\lesssim\inf_{\begin{subarray}{c}\xi_{\delta}\in\Lambda_{\delta}\end{subarray}\atop{B^{T}(\lambda-\xi_{\delta})\perp Z}}\|\lambda-\xi_{\delta}\|_{\Lambda}+h^{r}\|u-z\|_{W^{r}}.$
Moreover, if 3.2 holds, setting $u_{h}=A^{+}_{h}(f+B^{T}\lambda_{\delta})+z^{*}$ we have that
$\|u-u_{h}\|_{V}\lesssim\inf_{{\xi_{\delta}\in\Lambda_{\delta}}\atop{B^{T}(\lambda-\xi_{\delta})\perp Z}}\|\lambda-\xi_{\delta}\|_{\Lambda}+h^{r}\|u-z\|_{W^{r}}.$
Proof.
Let $\xi_{\delta}\in\Lambda_{\delta}$ be any element satisfying
(3.4) $b(z,\lambda-\xi_{\delta})=0.$
We have
$\|\lambda_{\delta}-\xi_{\delta}\|_{\Lambda}\lesssim s_{h}(\lambda_{\delta}-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})\lesssim s_{h}(\lambda-\xi_{\delta},\lambda_{\delta}-\xi_{\delta})$
$+[(s(\lambda,\lambda_{\delta}-\xi_{\delta})-\langle g,\lambda_{\delta}-\xi_{\delta}\rangle)-(s_{h}(\lambda,\lambda_{\delta}-\xi_{\delta})-\langle g_{h},\lambda_{\delta}-\xi_{\delta}\rangle)]$
which easily yields
$\|\lambda_{\delta}-\xi_{\delta}\|_{\Lambda}\lesssim\|\lambda-\xi_{\delta}\|_{\Lambda}+\sup_{\mu_{\delta}\in\Lambda_{\delta}}\frac{|[s(\lambda,\mu_{\delta})-\langle g,\mu_{\delta}\rangle]-[s_{h}(\lambda,\mu_{\delta})-\langle g_{h},\mu_{\delta}\rangle]|}{\|\mu_{\delta}\|_{\Lambda}}.$
We now need to estimate the consistency errors. We observe that
$s(\lambda,\mu_{\delta})-\langle g,\mu_{\delta}\rangle=b(A^{+}(B^{T}\lambda+f),\mu_{\delta})$
and
$s_{h}(\lambda,\mu_{\delta})-\langle g_{h},\mu_{\delta}\rangle=b(A_{h}^{+}(B^{T}\lambda+f),\mu_{\delta}).$
Then
$|[s(\lambda,\mu_{\delta})-\langle g,\mu_{\delta}\rangle]-[s_{h}(\lambda,\mu_{\delta})-\langle g_{h},\mu_{\delta}\rangle]|=|b((A^{+}-A^{+}_{h})(B^{T}\lambda+f),\mu_{\delta})|$
$\lesssim\|B(A^{+}-A^{+}_{h})(B^{T}\lambda+f)\|_{\Lambda^{\prime}}\|\mu_{\delta}\|_{\Lambda}.$
Assumption 3.3 yields
$|[s(\lambda,\mu_{\delta})-\langle g,\mu_{\delta}\rangle]-[s_{h}(\lambda,\mu_{\delta})-\langle g_{h},\mu_{\delta}\rangle]|\lesssim h^{r}\|A^{+}(B^{T}\lambda+f)\|_{W^{r}}.$
Let us now bound $\|z-z^{*}\|_{V}$. We have
$\|z-z^{*}\|_{V}\lesssim\sup_{\mu\in\Lambda_{0}}\frac{b(z-z^{*},\mu)}{\|\mu\|_{\Lambda}}.$
Now, as $\Lambda_{0}\subset\Lambda_{\delta}$ we have
$b(z-z^{*},\mu)=\langle g-g_{h},\mu\rangle-s(\lambda,\mu)+s_{h}(\lambda_{\delta},\mu)\\ =\langle g-g_{h},\mu\rangle-s(\lambda,\mu)+s_{h}(\lambda,\mu)+s_{h}(\lambda_{\delta}-\lambda,\mu)\\ \lesssim h^{r}\|A^{+}(B^{T}\lambda+f)\|_{W^{r}}\|\mu\|_{\Lambda}+\|\lambda_{\delta}-\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
Finally, let us estimate the error on $u$. We have
$u-u_{h}=A^{+}(f+B^{T}\lambda)-A^{+}_{h}(f+B^{T}\lambda_{\delta})+z-z^{*}\\ =(A^{+}-A^{+}_{h})(f+B^{T}\lambda)+A^{+}_{h}B^{T}(\lambda-\lambda_{\delta})+z-z^{*}.$
Therefore
$\|u-u_{h}\|_{V}=\|(A^{+}-A^{+}_{h})(f+B^{T}\lambda)+A^{+}_{h}B^{T}(\lambda-\lambda_{\delta})+z-z^{*}\|_{V}\\ \leq\underbrace{\|(A^{+}-A^{+}_{h})(f+B^{T}\lambda)\|_{V}}_{I}+\underbrace{\|A^{+}_{h}B^{T}(\lambda-\lambda_{\delta})\|_{V}}_{II}+\|z-z^{*}\|_{V}\\ \lesssim h^{r}\|u-z\|_{W^{r}}+\|\lambda-\lambda_{\delta}\|_{\Lambda}+\|z-z^{*}\|_{V},$
where we used Assumption 3.2 and (2.6) for $I$ followed by the boundedness of $A^{+}_{h}$ and the bound $\|B^{T}(\lambda-\lambda_{\delta})\|_{V^{\prime}}\lesssim\|\lambda-\lambda_{\delta}\|_{\Lambda}$ for $II$. ∎
It remains to show that $\lambda$ can be optimally approximated by functions $\xi_{\delta}$ in $\Lambda_{\delta}$ with $B^{T}(\lambda-\xi_{\delta})\perp Z$. This is the object of the following proposition.
Proposition 3.9.
Under Assumption 3.7 it holds that
(3.5) $\inf_{{\xi_{\delta}\in\Lambda_{\delta}}\atop{B^{T}(\lambda-\xi_{\delta})\perp Z}}\|\lambda-\xi_{\delta}\|_{\Lambda}\lesssim\inf_{{\xi_{\delta}\in\Lambda_{\delta}}}\|\lambda-\xi_{\delta}\|_{\Lambda}$
Proof.
It is sufficient to take $\xi_{\delta}$ as the solution to the problem
$s(\xi_{\delta},\mu_{\delta})+b(z^{\prime},\mu_{\delta})$ $=$ $s(\lambda,\mu_{\delta})\qquad\forall\mu_{\delta}\in\Lambda_{\delta}$
$b(w,\xi_{\delta})$ $=$ $b(w,\lambda),\qquad\forall w\in Z.$
Standard error estimates for the solution of saddle point problems yield (3.5). ∎
Corollary 3.10.
Under Assumptions 3.2, 3.4, 3.5 and 3.7, if $u\in W^{r}$, $r\leq R$ and $\lambda\in\Xi^{t}$, $t\leq m$, setting $u_{h}=A^{+}_{h}(f+B^{T}\lambda_{\delta})+z^{*}$ we have that
$\|u-u_{h}\|_{V}\lesssim\delta^{t}\|\lambda\|_{\Xi^{t}}+h^{r}\|u-z\|_{W^{r}}.$
Remark 3.11.
Note that it is not the smoothness of $u$ that comes into play in Theorem 3.8, but the smoothness of $u-z$.
3.1. Including stabilization
Consider now a situation where the condition $h/\delta<C_{\Lambda}$ with $C_{\Lambda}$ sufficiently small, given by Lemma 3.5, does not hold, but where we can not (or do not want to) take a smaller $h$ or a larger $\delta$. Assume however that we know that (3.2) holds, with a coarser mesh-size $\tilde{\delta}$ for which $h/\tilde{\delta}<C_{\Lambda}$, for some auxiliary space $\widetilde{\Lambda}_{\delta}$, not necessarily related to $\Lambda_{\delta}$, but with similar asymptotic properties of approximation. This can be exploited to design a stabilization term and retrieve the well posedness of the discrete problem.
To this aim we let $\tilde{\pi}_{\delta}:\Lambda\to\tilde{\Lambda}_{\delta}$ denote a bounded projection on $\tilde{\Lambda}_{\delta}$. Moreover we let $[\cdot,\cdot]_{\delta}:(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\times(\Lambda_{\delta}+\tilde{\Lambda}_{\delta})\to\mathbb{R}$ denote a continuous bilinear form satisfying
$[\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}]_{\delta}\gtrsim\|\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta}\|^{2}_{\Lambda},\qquad\text{ for all }\lambda_{\delta}\in\Lambda_{\delta},$
and we set
$j(\lambda_{\delta},\mu_{\delta})=[\lambda_{\delta}-\tilde{\pi}_{\delta}\lambda_{\delta},\mu_{\delta}-\tilde{\pi}_{\delta}\mu_{\delta}]_{\delta}.$
We consider the following stabilized version of Problem 3.1, where $\gamma\in\mathbb{R}^{+}$ is a stabilization parameter the size of which will influence the stability of the system.
Problem 3.12.
Find $\lambda_{\delta}\in\Lambda_{\delta}$ and $z^{*}\in Z$ such that
$s_{h}(\lambda_{\delta},\mu_{\delta})+\gamma j(\lambda_{\delta},\mu_{\delta})+b(z^{*},\mu_{\delta})$ $=$ $\langle g_{h},\mu_{\delta}\rangle,\qquad\forall\mu_{\delta}\in\Lambda_{\delta},$
$b(w,\lambda_{\delta})$ $=$ $-\langle f,w\rangle\qquad\forall w\in Z.$
We have the following theorem.
Theorem 3.13.
Assume that the black box approximate solver $A_{h}^{+}$ satisfies Assumptions 3.3 and 3.7, and that the space $\tilde{\Lambda}_{\delta}$ satisfies Assumption 3.5. Then, there exists $\gamma_{0}$ such that, provided $\gamma>\gamma_{0}$, Problem 3.12 is well posed, and the following error bound holds for its solution:
$\|\lambda-\lambda_{\delta}\|_{\Lambda}+\|z-z^{*}\|_{V}\lesssim\inf_{{\xi_{\delta}\in\Lambda_{\delta}}\atop{B^{T}(\lambda-\xi_{\delta})\perp Z}}\|\lambda-\xi_{\delta}\|_{\Lambda}+\inf_{\xi_{\tilde{\delta}}\in\tilde{\Lambda}_{\delta}}\|\lambda-\xi_{\tilde{\delta}}\|_{\Lambda}+h^{r}\|u-z\|_{W^{r}}.$
Moreover, if Assumption 3.2 holds, setting $u_{h}=A^{+}_{h}(f+B^{T}\lambda_{\delta})+z^{*}$ we have that
$\|u-u_{h}\|_{V}\lesssim\inf_{{\xi_{\delta}\in\Lambda_{\delta}}\atop{B^{T}(\lambda-\xi_{\delta})\perp Z}}\|\lambda-\xi_{\delta}\|_{\Lambda}+\inf_{\xi_{\tilde{\delta}}\in\tilde{\Lambda}_{\delta}}\|\lambda-\xi_{\tilde{\delta}}\|_{\Lambda}+h^{r}\|u-z\|_{W^{r}}.$
Proof.
To prove well posedness we follow the proof of Lemma 3.6. For $\lambda_{\delta}\in\Lambda_{\delta}\cap\widehat{\Lambda}$ we can write
[…]
In the first term we first add and subtract $-BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta}$ on the left slot to get
[…]
and then, letting $c_{s}$ be the coercivity constant of $s(\cdot,\cdot)$, we bound from below the second term on the right hand side by adding and subtracting $-BA^{+}B^{T}\lambda_{\delta}$, which allows us to write
$BA^{+}B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>=\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+s(\lambda_{\delta},\lambda_{\delta})\\ \geq\left<BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>+c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
Now we can bound
$BA^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq C\|\tilde{\pi}_{\delta}\lambda_{\delta}-\lambda_{\delta}\|_{\Lambda}\|\lambda_{\delta}\|_{\Lambda}\leq C\sqrt{j(\lambda_{\delta},\lambda_{\delta})}\|\lambda_{\delta}\|_{\Lambda}\\ \leq c_{1}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
Analogously we have that
$BA_{h}^{+}B^{T}(\tilde{\pi}_{\delta}-I)\lambda_{\delta},\lambda_{\delta}\right>\leq c_{2}\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta})+\frac{1}{4}\epsilon\|\lambda_{\delta}\|_{\Lambda}^{2}.$
Using now the assumed properties of the black box approximate solver $A^{+}_{h}$ we have that
$B(A_{h}^{+}-A)B^{T}\tilde{\pi}_{\delta}\lambda_{\delta},\lambda_{\delta}\right>\leq Ch^{r}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{\Xi^{r}}\|\lambda_{\delta}\|_{\Lambda}\leq\frac{h^{r}}{\tilde{\delta}^{r}}\|\tilde{\pi}_{\delta}\lambda_{\delta}\|_{V}\|\lambda_{\delta}\|_{\Lambda}\leq c_{3}\frac{h^{r}}{\tilde{\delta}^{r}}\|\lambda_{\delta}\|_{\Lambda}^{2}.$
Collecting terms we see that
$s_{h}(\lambda_{\delta},\lambda_{\delta})\geq\left(c_{s}-\frac{1}{2}\epsilon-c_{3}\left(\frac{h}{\tilde{\delta}}\right)^{r}\right)\|\lambda_{\delta}\|_{\Lambda}^{2}-2(c_{1}+c_{2})\epsilon^{-1}j(\lambda_{\delta},\lambda_{\delta}).$
Assuming that $\epsilon=1/2c_{s}$ and
$c_{3}\left(\frac{h}{\tilde{\delta}}\right)^{r}\leq\frac{1}{4}c_{s}$
the partial coercivity follows
(3.6) $s_{h}(\lambda_{\delta},\lambda_{\delta})\geq\frac{1}{2}c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}-4c_{s}(c_{1}+c_{2})j(\lambda_{\delta},\lambda_{\delta}).$
Considering this bound in the stability estimate of Problem 3.12 we see that
$s_{h}(\lambda_{\delta},\lambda_{\delta})+\gamma j(\lambda_{\delta},\lambda_{\delta})\geq\frac{1}{2}c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}+(\gamma-4c_{s}(c_{1}+c_{2}))j(\lambda_{\delta},\lambda_{\delta})$
and hence the stabilized method has similar stability properties as the inf-sup stable method if $\gamma\geq 4c_{s}(c_{1}+c_{2})$.
We conclude that for such a choice of $\gamma$,
(3.7) $\frac{1}{2}c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}\leq s_{h}(\lambda_{\delta},\lambda_{\delta})+\gamma j(\lambda_{\delta},\lambda_{\delta}).$
For the error analysis we proceed as in Theorem 3.8, using now the bound (3.7). Let once again $\xi_{\delta}\in\Lambda_{\delta}$ satisfy (3.4), and let $e_{h}=\lambda_{\delta}-\xi_{\delta}$, with now $\lambda_{\delta}$ solution to Problem 3.12. We have
$\frac{1}{2}c_{s}\|e_{h}\|_{\Lambda}^{2}\leq s_{h}(e_{h},e_{h})+\gamma j(e_{h},e_{h}).$
We see that
$s_{h}(e_{h},e_{h})+\gamma j(e_{h},e_{h})=s_{h}(\lambda-\xi_{\delta},e_{h})+\langle g_{h}-g,e_{h}\rangle+s(\lambda,e_{h})-s_{h}(\lambda,e_{h})-\gamma j(\xi_{\delta},e_{h}).$
The first four terms are bounded as in the proof of Theorem 3.8. As far as the last term is concerned, we observe that
$j(\xi_{\delta},e_{h})\lesssim\|\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta}\|_{\Lambda}\|e_{h}\|_{\Lambda}.$
Now, adding and subtracting $(I-\tilde{\pi}_{\delta})\lambda$ we see that
$\|\xi_{\delta}-\tilde{\pi}_{\delta}\lambda\|_{\Lambda}\leq\|\lambda-\tilde{\pi}_{\delta}\lambda\|_{\Lambda}+\|(I-\tilde{\pi}_{\delta})(\lambda-\xi_{\delta})\|_{\Lambda}\lesssim\|\lambda-\tilde{\pi}_{\delta}\lambda\|_{\Lambda}+\|\lambda-\xi_{\delta}\|_{\Lambda}.$
Then, proceeding as in the proof of Theorem 3.8, we obtain the bound on $\|\lambda-\lambda_{\delta}\|_{\Lambda}$. The bounds on $\|z-z^{*}\|_{V}$ and on $\|u-u_{h}\|_{V}$ also follow using the argument of Theorem 3.8. ∎
The lower bound $\gamma>\gamma_{0}$ on the stabilization parameter, required by Theorem 3.13, can be relaxed if $A_{h}^{+}$ is a positive semi definite operator, so that $s_{h}(\lambda_{\delta},\lambda_{\delta})\geq 0$ for all $\lambda_{\delta}\in\Lambda_{\delta}$. More precisely, we have the following corollary.
Corollary 3.14.
Under the assumptions of Theorem 3.13, if $s_{h}$ is positive semi-definite then the error bounds hold for all $\gamma>0$, with a hidden constant that scales as $\gamma^{-1}$.
Proof.
We only need to show that coercivity holds for all $\gamma>0$. To this end we observe that, as $s_{h}(\lambda_{\delta},\lambda_{\delta})\geq 0$, for $\eta\in(0,1)$ arbitrary we have that
$s_{h}(\lambda_{\delta},\lambda_{\delta})+\gamma j(\lambda_{\delta},\lambda_{\delta})\geq\eta s_{h}(\lambda_{\delta},\lambda_{\delta})+\gamma j(\lambda_{\delta},\lambda_{\delta}).$
Using (3.6) to bound the first term of the right hand side we get
$s_{h}(\lambda_{\delta},\lambda_{\delta})+\gamma j(\lambda_{\delta},\lambda_{\delta})\geq\frac{\eta}{2}c_{s}\|\lambda_{\delta}\|_{\Lambda}^{2}+(\gamma-4\eta c_{s}(c_{1}+c_{2}))j(\lambda_{\delta},\lambda_{\delta}).$
We obtain coercivity by choosing $\eta<\gamma/(4c_{s}((c_{1}+c_{2}))$ and the claim follows. ∎
Corollary 3.14 applies, in particular, to the case where, mimicking (2.5) on the discrete level, for all $k$, $A_{h,k}^{+}$ is defined through a Galerkin projection on a finite dimensional subspace $C_{h,k}\subset C_{k}$. Indeed, for $g\in V_{k}^{\prime}$, we can define $A^{+}_{h,k}g\in C_{h,k}$ as the solution to
(3.8) $a_{k}(A^{+}_{h,k}g,w_{h})=\langle g,w_{h}\rangle,\qquad\text{ for all }w_{h}\in C_{h,k}.$
Then, to prove the positive semi definiteness of $s_{h}$, using (3.8) and the coercivity of $a_{k}$ on $C_{k}$ we can write, for all $\lambda_{\delta}\in\Lambda_{\delta}$,
$0\leq a_{k}(A^{+}_{h,k}B^{T}\lambda_{\delta},A^{+}_{h,k}B^{T}\lambda_{\delta})=\langle B^{T}\lambda_{\delta},A^{+}_{h,k}B^{T}\lambda_{\delta}\rangle\\ =\langle\lambda_{\delta},BA^{+}_{h,k}B^{T}\lambda_{\delta}\rangle=s_{h}(\lambda_{\delta}.\lambda_{\delta}).$
Remark 3.15.
We note that since $\Lambda_{\delta}$ is finite dimensional, it is often possible to choose $j(\xi_{\delta},\xi_{\delta})$ of a form convenient for computation. Indeed in some cases the projection operator can be eliminated and replaced by some other operator acting directly on $\xi_{\delta}$. This is the case for instance when, in the domain decomposition framework (see Section 5.1), $\widetilde{\Lambda}_{\delta}$ and $\Lambda_{\delta}$ are different finite element spaces of different mesh size, but with the same polynomial order. Then typically $[\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta},\xi_{\delta}-\tilde{\pi}_{\delta}\xi_{\delta}]_{\delta}$ can be replaced by a penalty term acting on the jumps of derivatives of $\xi_{\delta}$ alone, see [14]. In such a framework, computable bilinear forms $[\cdot,\cdot]_{\delta}$ can also be designed by resorting to suitable localization results, see [8]. For an alternative way of stabilizing in an abstract framework, see also [5].
4. Solving the resulting discrete problem
Since we are interested in large problems, it is natural to consider an iterative solver for the reduced system of Problem 3.1, where the stiffness matrix relative to the approximate Schur complement operator need not be assembled, but where, at each iteration, the approximate black box solvers are called upon to evaluate their action on a given element of $\Lambda_{\delta}$. We note that for very large systems it may be advantageous to proceed in a nested fashion, where also the subproblems are themselves decomposed in smaller subproblems, until the local solvers are so small that direct solvers can be applied. Focusing on the unstabilized Problem 3.1, and following once again an approach inspired by the FETI method, in order to efficiently solve Problem 3.1 we will need
  1. a)
    to reduce the saddle point problem posed on $\Lambda_{\delta}\times Z$ to a positive definite problem defined on the subspace $\widehat{\Lambda}_{\delta}$ of $\Lambda_{\delta}$, which can be solved by a Krylov type method (e.g. PCG, if the corresponding operator is symmetric)
  2. b)
    to construct a preconditioner for the the restriction of the bilinear form $s_{h}$ to the subspace $\widehat{\Lambda}_{\delta}$.
The space $\widehat{\Lambda}_{\delta}$ and its dual
We will need, in the following, to simultaneously handle the dual spaces $\Lambda_{\delta}^{\prime}$ and $\widehat{\Lambda}_{\delta}^{\prime}$ of, respectively, $\Lambda_{\delta}$ and $\widehat{\Lambda}_{\delta}=\Lambda_{\delta}\cap\ker G^{T}$, and, in doing this, we need to take into account the relation between these two duals. To this aim, it will be useful to think of $\widehat{\Lambda}_{\delta}$ as an independent space and, to make this point of view explicit, we introduce the natural injection operator $\mathcal{I}:\widehat{\Lambda}_{\delta}\to\Lambda_{\delta}$. The space $\widehat{\Lambda}_{\delta}^{\prime}$ is isomorphic to the quotient space $\Lambda_{\delta}^{\prime}/\ker\mathcal{R}$ where $\mathcal{R}=\mathcal{I}^{T}:\Lambda_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}^{\prime}$ is the natural restriction operator, defined as
$\langle\mathcal{R}\phi,\widehat{\lambda}\rangle=\langle\phi,\mathcal{I}\widehat{\lambda}\rangle=\langle\phi,\widehat{\lambda}\rangle.$
The elements of $\widehat{\Lambda}_{\delta}^{\prime}$ are then naturally defined as equivalence classes ${\widetilde{\varphi}}=[\phi]$, $\phi\in\Lambda_{\delta}^{\prime}$, where $\phi^{\prime}\in[\phi]$ iff $\mathcal{R}(\phi-\phi^{\prime})=0$. In constructing our solver, we will need to identify elements of $\widehat{\Lambda}_{\delta}^{\prime}$ with a canonical representative $\phi\in\Lambda_{\delta}^{\prime}$ minimizing a suitable norm. To this aim we introduce a discrete scalar product $\sigma:\Lambda_{\delta}\times\Lambda_{\delta}\to\mathbb{R}$, not necessarily equivalent to the $\Lambda$ scalar product, and we define a mapping $\mathcal{K}:\widehat{\Lambda}_{\delta}^{\prime}\to\Lambda_{\delta}^{\prime}$ with, for all ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$, ${\widetilde{\varphi}}=[\mathcal{K}{\widetilde{\varphi}}]$, defined as
$\mathcal{K}{\widetilde{\varphi}}=\mathop{\text{argmin}}_{{\phi^{\prime}\in\Lambda_{\delta}^{\prime}}\ :\ {{\widetilde{\varphi}}=[\phi^{\prime}]}}\|\phi^{\prime}\|_{\sigma,*},\qquad\text{ where }\quad\|\phi^{\prime}\|_{\sigma,*}=\sup_{\lambda\in\Lambda_{\delta}}\frac{\langle\phi^{\prime},\lambda\rangle}{\|\lambda\|_{\sigma}}.$
We have the following proposition clarifying the relation between the canonical representative $\mathcal{K}({\widetilde{\varphi}})$ of an element ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$ and any element $\phi\in\Lambda_{\delta}^{\prime}$ such that $[\phi]={\widetilde{\varphi}}$.
Proposition 4.1.
Let $\Pi_{\sigma}:\Lambda_{\delta}\to\widehat{\Lambda}_{\delta}$ denote the $\sigma$-orthogonal projector onto $\widehat{\Lambda}_{\delta}$. Then for all $\phi\in\Lambda_{\delta}^{\prime}$ it holds
$\mathcal{K}([\phi])=\Pi_{\sigma}^{T}\phi.$
Proof.
It is easy to check that $\Pi_{\sigma}^{T}\phi$ verifies $\Pi_{\sigma}^{T}\phi\in[\phi]$. Indeed, for $\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$ we have
$\langle\phi-\Pi_{\sigma}^{T}\phi,\widehat{\lambda}\rangle=\langle\phi,\widehat{\lambda}-\Pi_{\sigma}\widehat{\lambda}\rangle=0.$
Let now $\phi^{\prime}\in[\phi]$. We have
$\langle\Pi_{\sigma}^{T}\phi^{\prime}-\Pi_{\sigma}^{T}\phi,\lambda\rangle=\langle\phi^{\prime}-\phi,\Pi_{\sigma}\lambda\rangle=0.$
Then
$\|\Pi_{\sigma}^{T}\phi\|_{\sigma,*}=\|\Pi_{\sigma}^{T}\phi^{\prime}\|_{\sigma,*}=\sup_{\lambda\in\Lambda_{\delta}}\frac{\langle\Pi_{\sigma}^{T}\phi^{\prime},\lambda\rangle}{\|\lambda\|_{\sigma}}=\sup_{\lambda\in\Lambda_{\delta}}\frac{\langle\phi^{\prime},\Pi_{\sigma}\lambda\rangle}{\|\lambda\|_{\sigma}}\leq\|\phi^{\prime}\|_{\sigma,*}\sup_{\lambda\in\Lambda_{\delta}}\frac{\|\Pi_{\sigma}\lambda\|_{\sigma}}{\|\lambda\|_{\sigma}}\leq\|\phi^{\prime}\|_{\sigma,*},$
whence
$\|\Pi_{\sigma}^{T}\phi\|_{\sigma,*}=\min_{\phi^{\prime}\in[\phi]}\|\phi^{\prime}\|_{\sigma,*},$
which implies that $\mathcal{K}[\phi]=\Pi_{\sigma}^{T}\phi$. ∎
Observe that $\Pi_{\sigma}$ can be defined as $\Pi_{\sigma}\lambda=\eta\in\Lambda_{\delta}$ with $\eta$ being the first component of the solution $(\eta,\widetilde{z})\in\Lambda_{\delta}\times Z$ of the system
$\begin{array}[]{lcl}\sigma(\eta,\mu)+b(\widetilde{z},\mu)&=&\sigma(\lambda,\mu),\qquad\forall\mu\in\Lambda_{\delta}\\ b(z^{\prime},\eta)&=&0,\qquad\forall z^{\prime}\in Z.\end{array}$
Reduction to $\widehat{\Lambda}_{\delta}$
Letting $\lambda$ be the solution to problem 3.1, we have that $\lambda^{0}=\lambda-\Pi_{\sigma}\lambda$ is the first component of the solution $(\lambda^{0},z^{0})\in\Lambda_{\delta}\times Z$ to
(4.1) $\begin{array}[]{lcl}\sigma(\lambda^{0},\mu)-b(z^{0},\mu)&=&0,\qquad\forall\mu\in\Lambda_{\delta}\\ b(z^{\prime},\lambda^{0})&=&-\langle f,z^{\prime}\rangle,\qquad\forall z^{\prime}\in Z.\end{array}$
We assume that the scalar product $\sigma$ is chosen in such a way that such a problem can be solved cheaply (at least with respect to the cost of the solution of the whole Problem 3.1), and $\widehat{\lambda}=\Pi_{\sigma}\lambda=\lambda-\lambda^{0}$ is the solution to the following reduced problem
(4.2) $s_{h}(\widehat{\lambda},\widehat{\mu})=\langle g,\widehat{\mu}\rangle-\langle S_{h}\lambda^{0},\widehat{\mu}\rangle,\qquad\forall\widehat{\mu}\in\widehat{\Lambda}_{\delta}.$
We know by Lemma 3.6 that $s_{h}$ is coercive on $\widehat{\Lambda}_{\delta}$, and, consequently, that the corresponding operator $\widehat{S}_{h}:\widehat{\Lambda}_{\delta}\to\widehat{\Lambda}_{\delta}^{\prime}$ is invertible. The problem can be then addressed by any iterative solver well suited to solve linear systems with positive definite matrices. For instance, if, in addition, the operator $A^{+}_{h}$ is symmetric, which implies that $S_{h}$ and, consequently, $\widehat{S}_{h}$ are symmetric, one can resort to the preconditioned conjugate gradient method.
Construction of the preconditioner $\widehat{M}$
What we wold like, in order to build a preconditioner, is something resembling an inverse of $S_{h}$, if $S_{h}$ was invertible. If $A^{+}_{h}$ and $B_{\delta}$ were also, in turn, invertible, this would be
(4.3) $(BA^{+}_{h}B^{T})^{-1}=B^{-T}(A^{+}_{h})^{-1}B^{-1}.$
Of course, none of the above mentioned operators is generally invertible; the idea is then to use the right-hand side expression as a guideline, and replace $B^{-1}$ and $(A_{h}^{+})^{-1}$, respectively, with a suitable pseudo inverse $B_{\delta}^{+}$, and with $A$ (or some spectrally equivalent $\widetilde{A}$).
Following the approach put forward by the FETI preconditioner, we rely on three ingredients The first ingredient is a finite dimensional space $V_{h}\subseteq V$, which we assume to satisfy
(4.4) $Z\subseteq V_{h},\qquad\text{and}\qquad\inf_{\lambda\in\Lambda_{\delta}}\sup_{v\in V_{h}}\frac{b(v,\lambda)}{\|v\|_{V}\|\lambda\|_{\Lambda}}\gtrsim 1.$
Observe that, if a space $C_{h}$ underlying the approximate pseudo inverse $A^{+}_{h}$ is available, a natural choice for $V_{h}$ is $V_{h}=C_{h}$, but we are also interested in situations in which such a space might not be accessible. Also note that $V_{h}$ needs not satisfy any approximation assumption.
We let $A_{\delta}:V_{h}\to V_{h}^{\prime}$ and $B_{\delta}:V_{h}\to\Lambda_{\delta}^{\prime}$, be the discrete versions of the operators $A$ and $B$, defined as
$\langle A_{\delta}v,w\rangle=a(v,w)\ \forall v,w\in V_{h},\qquad\langle B_{\delta}v,\lambda\rangle=b(v,\lambda)\ \forall v\in V_{h},\lambda\in\Lambda_{\delta}.$
The second ingredient is a discrete scalar product $d:V_{h}\times V_{h}\to\mathbb{R}$ (not necessarily equivalent to the natural scalar product for $V$) and a corresponding $d$-orthogonal splitting of the space $V_{h}$ as
$V_{h}=\ker B_{\delta}\oplus(\ker B_{\delta})^{\perp},$
where
$\ker B_{\delta}=\{v\in V_{h}:\ b(v,\mu)=0,\ \forall\mu\in\Lambda_{\delta}\},$
$(\ker B_{\delta})^{\perp}=\{v\in V_{h}:\ d(v,w)=0,\ \forall w\in\ker B_{\delta}\}.$
We let $\Pi_{d}:V_{h}\to(\ker B_{\delta})^{\perp}$ denote the $d$-orthogonal projection onto $(\ker B_{\delta})^{\perp}$. Whichever the choice of the scalar product $d$, the restriction of $B_{\delta}$ to $(\ker B_{\delta})^{\perp}$ is bijective, and we let $B_{\delta}^{+}:\Lambda_{\delta}^{\prime}\to(\ker B_{\delta})^{\perp}$ denote its inverse, which is a pseudo inverse of $B_{\delta}$, verifying
(4.5) $B_{\delta}B_{\delta}^{+}=\mathbf{1}_{{\Lambda_{\delta}^{\prime}}},\qquad B_{\delta}^{+}B_{\delta}=\Pi_{d}.$
We easily realize that, given $\phi\in\Lambda_{\delta}^{\prime}$, the pseudo inverse $B_{\delta}^{+}\phi$ can be defined as $B_{\delta}^{+}\phi=v$, where $(v,\zeta)\in V_{h}\times\Lambda_{\delta}$ is the solution to
$\begin{array}[]{lcl}d(v,w)+b(w,\zeta)&=&0,\qquad\forall w\in V_{h}\\ b(v,\mu)&=&\langle\phi,\mu\rangle,\forall\mu\in\Lambda_{\delta}.\end{array}$
Following the guidelines provided by the formal inversion expression (4.3) we introduce the linear operator $M:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}$, defined as
$M=(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+},$
which corresponds to the bilinear form $m:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}$ defined as
$m(\phi,\psi)=a(B_{\delta}^{+}\phi,B_{\delta}^{+}\psi).$
We finally define our preconditioner $\widehat{M}:\widehat{\Lambda}_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}$ as the operator corresponding to the bilinear form $\widehat{m}:\widehat{\Lambda}_{\delta}^{\prime}\times\widehat{\Lambda}_{\delta}^{\prime}\to\mathbb{R}$, defined as
$\widehat{m}({\widetilde{\varphi}},\widetilde{\psi})=m(\mathcal{K}{\widetilde{\varphi}},\mathcal{K}\widetilde{\psi}).$
We have the following theorem, the proof of which is fundamentally the same as for the FETI method in domain decomposition framework (see [29]). For the sake of completeness we report it in Appendix B.
Theorem 4.2.
Let $\Pi_{d}:V_{h}\to V_{h}$ denote the ${d}$-orthogonal projection operator onto the ${d}$-orthogonal complement $(\ker B_{\delta})^{\perp}$ of $\ker B_{\delta}$ in $V_{h}$. Then it holds
${\kappa(\widehat{M}\widehat{S}_{h})\lesssim C(V_{h},d,\sigma)}$
with
$C(V_{h},d,\sigma)=\|\Pi_{d}\|_{V\to V}\left(1+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}\|_{V\to V}\right).$
4.1. Strong form of the iterative solver
Let us give some detail on how the different steps involved in the reduction of the discrete problem and in the application of the preconditioner can be split as composition of a small numbers of linear operators, which are the ones that need to be implemented. To this aim we introduce the Riesz operators relative to the two scalar products $\sigma$ and $d$, namely $\Sigma:\Lambda_{\delta}\to\Lambda_{\delta}^{\prime}$ and $D:V_{h}\to V_{h}^{\prime}$ be defined as
$\langle\Sigma\lambda,\mu\rangle=\sigma(\lambda,\mu),\ \forall\lambda,\mu\in\Lambda_{\delta},\qquad\langle Dv,w\rangle=d(v,w),\ \forall v,w\in V_{h}.$
Moreover we let $G_{\delta}:Z\to\Lambda_{\delta}^{\prime}$ denote the restriction to $Z$ of the operator $B_{\delta}$:
$\langle G_{\delta}z,\lambda\rangle=\langle B_{\delta}z,\lambda\rangle=b(z,\lambda),\ \forall z\in Z,\ \lambda\in\Lambda_{\delta}.$
Remark that we need to distinguish between the operators $B_{\delta}$ and its restriction $G_{\delta}$ as we will need to take the adjoint operators. Some standard linear algebra manipulations allow to see that the projectors $\Pi_{d}$ and $\Pi_{\sigma}$ verify respectively
$\Pi_{d}=D^{-1}B_{\delta}^{T}(B_{\delta}D^{-1}B_{\delta}^{T})^{-1}B_{\delta},\qquad\Pi_{\sigma}=1-\Sigma^{-1}G_{\delta}(G_{\delta}^{T}\Sigma^{-1}G_{\delta})^{-1}G_{\delta}^{T}.$
while the reduced equation (4.2) in strong form reads
(4.6) $\mathcal{I}^{T}S_{h}\mathcal{I}\eta=\mathcal{I}^{T}g-\mathcal{I}^{T}S_{h}\lambda^{0},$
with
$\lambda^{0}=-\Sigma^{-1}G(G^{T}\Sigma^{-1}G)^{-1}f.$
The pseudo inverse $B_{\delta}$ can be realized as
$B_{\delta}^{+}=D^{-1}B_{\delta}^{T}(B_{\delta}D^{-1}B_{\delta}^{T})^{-1},$
and the linear operator corresponding to the bilinear form $m$ is
$M=(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+}.$
Finally, exploiting Proposition 4.1 we can write
$\widehat{M}=\Pi_{\sigma}(B_{\delta}^{+})^{T}A_{\delta}B_{\delta}^{+}\Pi_{\sigma}^{T}.$
Remark that, while the operators $A_{\delta}$, $B_{\delta}$ and $G_{\delta}$ are substantially given by the continuous problem, there is, a priori, a certain freedom in the choice of the operators $D$ and $\Sigma$ (or, equivalently, of the scalar products $d$ for $V_{h}$ and $\sigma$ for $\Lambda_{\delta}$). These have to be chosen in such a way that the resulting preconditioner $\widehat{M}$ is cheap (or easily parallelizable), and that, at the same time, the norms of $\Pi_{d}$ and $B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}$ stay bounded. How this can be done depends on the actual problem considered. Remark that choosing $\sigma$ and $d$ as the natural scalar products for $\Lambda$ and $V$ would immediately result in the boundedness of the above mentioned operators but would possibly lead to expensive preconditioner.
5. Applications
The above framework has multiple possible applications. The most obvious one is of course to design domain decomposition solvers for large scale problems. However it can also be used for the design of multiphysics, multiscale or mixed-dimensional problems. Here we show some typical model situations that enter the framework. We develop the domain decomposition case in the general setting of a multi-domain coupling. In the other examples, for the sake of notational simplicity, we consider model situations where only two or three domains are coupled. The extension to the multi-domain case is straightforward using the arguments of the first example.
5.1. Domain decomposition with Neumann or Dirichlet coupling
Let $\Omega\subset\mathbb{R}^{2}$ be a polygonal domain with boundary $\Gamma$. We consider the following simple model problem: given $f\in L^{2}(\Omega)$, find $u$ satisfying
[…]
where the diffusion coefficient $\kappa$ satisfies $0<\kappa_{0}\leq\kappa(x)\leq M$, with $\kappa_{0},M$ positive constants. We let $\Omega$ be split as the non-overlapping union of polygonal subdomains: $\overline{\Omega}=\cup_{k=1}^{K}\overline{\Omega}^{k}$ and set $\Gamma^{k}=\partial\Omega^{k}$. For each $k$ we let
$V_{k}=\{u\in H^{1}(\Omega^{k}),~{}u|_{\Gamma\cap\Gamma^{k}}=0\}.$
Letting $\nu^{k}$ denote the outer unit normal to the subdomain $\Omega^{k}$, on the skeleton $\Sigma$ we next select a normal direction $\nu$, taking care that on $\Gamma$ $\nu$ coincides with the outer normal to $\Omega$. We let $\Lambda$ denote the closure of $L^{2}(\Sigma)$ with respect to the norm
$\|\lambda\|_{-1/2,\Sigma}^{2}=\sum_{k}\|(\nu\cdot\nu^{k})\lambda\|_{-1/2,\Gamma_{k}}^{2}.$
The set $\Lambda$ is well suited to describe the trace on $\Sigma$ of the normal derivative of the solution to our problem. Equation (5.1) can be split as a system of couples local problems as: find $(u^{k})_{k}\in V=\prod_{k}V_{k}$, $\lambda\in\Lambda$, such that
$\int_{\Omega^{k}}\kappa\nabla u^{k}\cdot\nabla v^{k}-\int_{\Gamma^{k}}(\nu\cdot\nu^{k})\lambda v^{k}=\int_{\Omega^{k}}fv^{k},\qquad\text{ for all }v^{k}\in V_{k},\ \text{ for all }k,$
$\sum_{k}\int_{\Gamma^{k}}u^{k}(\nu\cdot\nu^{k})\mu=0,\ \text{ forall }\mu\in\Lambda.$
This is the problem from which the abstract framework took inspiration, and consequently it clearly falls within its purview.
A second possibility for splitting Problem 5.1 into coupled local problems is to introduce, as a multiplier, the trace of the solution on the skeleton of the decomposition as an independent variable. To this aim we then let $V_{k}=H^{1}(\Omega^{k})\times H^{-1/2}(\Gamma^{k})$ and the multiplier space is, this time, defined as $\Lambda=H^{1}(\Omega)|_{\Sigma}$. The coupled problem then reads: find $(u^{k},\phi^{k})_{k}\in\prod V_{k}$ and $\lambda\in\Lambda$ such that, for all $k$
$\Big{(}\int_{\Omega^{k}}\kappa\nabla u^{k}\cdot\nabla v^{k}-\int_{\Gamma^{k}}\phi^{k}v^{k}+\int_{\Gamma^{k}}u^{k}\psi^{k}\Big{)}-\int_{\Gamma^{k}}\lambda\psi^{k}=\int_{\Omega^{k}}fv^{k},\qquad\text{for all }(v^{k},\psi^{k})\in V_{k},$
and
$\sum_{k}\int_{\Gamma^{k}}\phi^{k}\mu=0,\qquad\text{ for all }\mu\in\Lambda.$
While this problem does not fully fit our abstract framework, as the bilinear form $a$ is not coercive but only inf-sup stable, the above problem is well posed and can be reduced to a well posed problem in the coupling variable $\lambda$, which can be handled as proposed. Observe that in the Dirichlet coupling there is no need to introduce norms with different orientations on the skeleton in order to define the coupling. On the discrete level any method for the weak imposition of Dirichlet conditions may be used. In particular, one may eliminate the local multiplier by applying Nitsche’s method in the spirit of [27]. For example if $\phi^{k}=\nu^{k}\cdot\kappa\nabla u^{k}$ and $\psi^{k}=\nu^{k}\cdot\kappa\nabla v^{k}$ a hybridised version of the penalty free Nitsche method introduced in [13] is obtained. Note however, that if the multiplier is eliminated, in general the coupling scheme can not be used in a black box fashion.
Remark 5.1.
Remark that, in the framework of this last domain decomposition formulation, Problem 3.1 substantially coincides with the natural domain decomposition method with non matching grids proposed in [25].
Remark 5.2.
We observe that, in the domain decomposition framework (but a similar thing might happen also in other situations), an inf-sup condition of the form (2.4) remains valid when neglecting the contribution from one of the two sides of the interface (that is, the supremum can be taken over the subspace of N-tuples $(v_{k})\in V$ with some of the components set to zero). At the discrete level, this implies that the stability of the bilinear form $s_{h}$ only requires Assumption 3.3 to hold only for the solvers corresponding to the non zero components. This is particularly useful in situations where different methods are coupled, where only some can readily be made to satisfy the conditions, whereas for the others it is less evident (for example when finite elements are coupled to a particle method).
5.2. Hybridized Johnson-Nedelec FEM - BEM coupling
We consider the Poisson equation, obtained by setting $\kappa=1$ in (5.1), in an exterior domain $\Omega=\mathbb{R}^{2}\setminus\Omega^{0}$, with $\Omega^{0}\subset\mathbb{R}^{2}$ bounded domain with boundary $\Gamma_{0}$. We assume that we have two disjoint connected subdomains $\Omega^{k}$, $k=1,2$, with $\Omega^{1}$ bounded and with $\bar{\Omega}^{0}\cap\bar{\Omega}^{2}=\varnothing$ , such that $\cup_{k=1}^{2}\bar{\Omega}^{k}\equiv\bar{\Omega}$, and we let $\Gamma:=\bar{\Omega}^{1}\cap\bar{\Omega}^{2}$. We assume that $f|_{\Omega^{1}}\in L^{2}(\Omega^{1})$, and that $f$ is compactly supported in $\Omega^{1}\cup\Gamma_{0}$. Problem (5.1) is well-posed in
$\mathscr{V}:=\{v:\omega(x)v(x)\in L^{2}(\Omega),\nabla v\in[L^{2}(\Omega)]^{2}\},$
where $\omega(x)=\left(\sqrt{1+|x|^{2}}\left(1+\log\sqrt{1+|x|^{2}}\right)\right)^{-1}$. We assume that $\Gamma_{0}$ and $\Gamma$ are smooth and we let $\nu^{k}$, $k=0,1$, be the normal to $\Gamma$, pointing outward of $\Omega^{k}$, and let $\gamma^{\Gamma}:H^{1}(\Omega)\mapsto H^{\frac{1}{2}}(\Gamma)$ denote the trace operator.
To write the problem as a system of two coupled subproblems, we consider the hybridized Johnson-Nedelec coupling introduced in [6]. To formulate the outer problem in $\Omega^{2}$ we introduce the single and double layer integral operators, $\mathcal{V}:H^{-\frac{1}{2}}(\Gamma)\mapsto H^{\frac{1}{2}}(\Gamma)$ and $\mathcal{K}:H^{\frac{1}{2}}(\Gamma)\mapsto H^{\frac{1}{2}}(\Gamma)$ defined by
$\mathcal{V}\psi(x):=\int_{\Gamma}G(x-y)\psi(s)~{}\mbox{d}\Gamma_{y},\qquad\text{and}\qquad\mathcal{K}\phi(x):=\int_{\Gamma}\partial_{\nu^{2}}G(x-y)\phi(s)~{}\mbox{d}\Gamma_{y},$
where $G(x):=-\tfrac{1}{2\pi}\log(|x|)$ is the fundamental solution of the Laplace equation. Using these operators it is well known that the exterior solution $u^{2}=u|_{\Omega^{2}}$ can be represented using the Kirchoff’s formula
(5.2) $u^{2}(x)=\mathcal{V}\psi-\mathcal{K}u^{2}(x)+\alpha,\quad x\in\Omega^{2},\alpha\in\mathbb{R}$
and
(5.3) $\frac{1}{2}u^{2}(x)=\mathcal{V}\psi-\mathcal{K}u^{2}(x)+\alpha,\quad x\in\Gamma,$
where $\psi:=\nu^{2}\cdot\nabla u^{2}|_{\Gamma}$, and where $\alpha\in\mathbb{R}$ is a constant related to the asymptotic behavior at infinity . To formulate the coupled problem we introduce the spaces
$V_{\Omega^{1}}=H^{1}_{0,\Gamma_{0}}(\Omega^{1}):=\{v\in H^{1}(\Omega^{1}):v|_{\Gamma^{0}}=0\},$
$V_{\Gamma}:=H^{\frac{1}{2}}(\Gamma),\quad V_{\Gamma}^{*}:=\{v\in V_{\Gamma}:\int_{\Gamma}v=0\},$
${\Lambda_{\Gamma}}:=H^{-\frac{1}{2}}(\Gamma),\quad\Lambda^{*}_{\Gamma}:=\{\mu\in H^{-\frac{1}{2}}(\Gamma):\int_{\Gamma}\mu=0\},$
and the bilinear forms $a_{1}:V_{\Omega^{1}}\times V_{\Omega^{1}}\to\mathbb{R}$, $a_{2,u}:\Lambda^{*}_{\Gamma}\times\Lambda^{*}_{\Gamma}\to\mathbb{R}$ and $a_{2,u}:V_{\Gamma}^{*}\times\Lambda^{*}_{\Gamma}\to\mathbb{R}$ defined as
$a_{1}(w,v):=\int_{\Omega^{1}}\nabla w\cdot\nabla v,\quad a_{2,\lambda}(\lambda,\mu):=2\langle\,\mathcal{V}\lambda,\mu\,\rangle_{\Gamma},\quad a_{2,u}(v,\mu):=-2\langle\,\mathcal{K}v,\mu\,\rangle_{\Gamma},$
and $b_{1}:V_{\Omega^{1}}\times{\Lambda_{\Gamma}}\to\mathbb{R}$ and $b_{2}:V_{\Gamma}\times{\Lambda_{\Gamma}}\to\mathbb{R}$, defined as
$b_{1}(v,\mu):=\langle\,\gamma_{\Gamma}v,\mu\,\rangle_{\Gamma},\qquad b_{2}(v,\mu):=-\langle\,v,\mu\,\rangle_{\Gamma},$
where $\langle\,\cdot,\cdot\,\rangle_{\Gamma}$ denotes the duality pairing between $H^{\frac{1}{2}}(\Gamma)$ and $H^{-\frac{1}{2}}(\Gamma)$. Letting $\lambda:=\nu^{1}\cdot\nabla u^{1}|_{\Gamma}$, the weak coupled formulation of the exterior problem then takes the form: find $u^{1},\lambda,u^{2},\psi\in V_{\Omega^{1}}\times{\Lambda_{\Gamma}}\times V_{\Gamma}\times\Lambda^{*}_{\Gamma}$ such that
(5.4) $a_{1}(u^{1},v^{1})-b_{1}(v^{1},\lambda)$ $=(f,v^{1})_{\Omega^{1}},\;\forall v^{1}\in V_{\Omega^{1}},$
(5.5) $a_{2,\lambda}(\psi,\tau)+a_{2,u}(u^{2},\tau)+b_{2}(u^{2},\tau)$ $=0,\;\forall\tau\in\Lambda^{*}_{\Gamma},$
(5.6) $-b_{2}(v^{2},\psi)-b_{2}(v^{2},\lambda)$ $=0,\;\forall v^{2}\in V_{\Gamma},$
(5.7) $b_{1}(u^{1},\mu)+b_{2}(u^{2},\mu)$ $=0\;\forall\mu\in{\Lambda_{\Gamma}}.$
For the well-posedness of this system we refer to [6]. In particular the inf-sup condition (2.4) is immediate, since $b_{i}$ is defined by the duality pairing between $H^{\frac{1}{2}}$ and $H^{-\frac{1}{2}}$ (see [6, Proposition 1]). In order to write such a problem in a way that can be framed in our abstract framework, we use (5.5) to obtain $\psi$ as a function of $u^{2}$, which, using (5.6) yields the equation
(5.8) $a_{2}(u^{2},v^{2}):=\langle\,\mathcal{V}^{-1}\left(\frac{1}{2}u^{2}+\mathcal{K}u^{2}\right),v^{2}\,\rangle=b_{2}(v^{2},\lambda),$
involving the Steklov-Poicaré operator $\mathcal{S}=\mathcal{V}^{-1}\left(\frac{1}{2}u^{2}+\mathcal{K}u^{2}\right)$, which replaces equations (5.5) and (5.6) in the coupled system. We can frame the resulting problem in our abstract framework by setting $V_{1}=V_{\Omega^{1}}$, $V_{2}=V_{\Gamma}$ and $\Lambda={\Lambda_{\Gamma}}$. Equations (5.4) and (5.8) play the role of equation (2.2) of the abstract framework, while equation (5.7) corresponds to equation (2.3). Due to the Dirichlet condition on $\Gamma^{0}$ the kernel of $A_{1}$, corresponding to the coercive bilinear form $a_{1}$, is trivial and $A^{+}_{1}$ is defined by the standard solution operator.
To show that $a_{2}$ is coercive on $V_{\Gamma}^{*}=(\ker A_{2})^{\perp}$, we let $\|v\|_{\mathcal{V}^{-1}}=\sqrt{\langle\mathcal{V}^{-1}v,v\rangle},$ and write $a_{2}(v,v)=\|v\|^{2}_{\mathcal{V}^{-1}}-\langle\,\mathcal{V}^{-1}\left(\frac{1}{2}v-\mathcal{K}v\right),v\,\rangle_{\Gamma}.$ Then we note that
$\langle\,\mathcal{V}^{-1}\left(\frac{1}{2}v-\mathcal{K}v\right),v\,\rangle_{\Gamma}\leq\|\frac{1}{2}v-\mathcal{K}v\|_{\mathcal{V}^{-1}}\|v\|_{\mathcal{V}^{-1}}\leq c_{K}\|v\|_{\mathcal{V}^{-1}}^{2}$
where $c_{K}<1$, see [26, Corollary 6.28]. Therefore $(1-c_{K})\|v\|_{\mathcal{V}^{-1}}^{2}\leq a_{2}(v,v)\quad\forall v\in V_{\Gamma}^{*}.$ It follows that the pseudo inverse $A_{2}^{+}$ is well defined on $V_{\Gamma}^{*}$ and the elements of the kernel $Z$ are the constant functions in the space $V_{\Gamma}$.
5.3. Coupling in the bulk or bulk to surface coupling
The framework can also be used in the nonstandard situation where the coupling takes place in a zone of the bulk where the two domains are overlapping. Such ideas have been applied to multiscale and multi-model couplings in [17, 2] but can also be used for overlapping domain decomposition using non-matching meshes. An advantage of the current framework is that the coupling field may be defined on a completely structured mesh in order to simplify the quadratures in the overlap. We propose two examples of problems of this type that can be modelled using the above framework.
5.3.1. Overlapping domain decomposition, Arlequin method
Here we consider a domain $\Omega$ with boundary $\Gamma$ and, once again, we wish to solve the Poisson equation (that is (5.1) with $\kappa=1$) in $\Omega$ with homogeneous boundary conditions on $\Gamma$. This time we decompose $\Omega$ into two overlapping subdomains $\Omega^{1}$ and $\Omega^{2}$ such that $\Omega=\Omega^{1}\cup\Omega^{2}$ and we let $\Omega^{\epsilon}=\Omega^{1}\cap\Omega^{2}$ denote some overlapping region of width $\epsilon$. Without loss of generality, we assume that $\partial\Omega^{1}\cap\partial\Omega$ has positive measure, while we do not assume that the same holds for $\partial\Omega^{2}\cap\partial\Omega$. Define $V^{k}:=\{v\in H^{1}(\Omega^{k}):v|_{{\Gamma^{D}}\cap\bar{\Omega}^{k}}=0\}$. Now we wish to solve independently on $\Omega^{1}$ and $\Omega^{2}$, and we can either couple at the interfaces of $\Omega^{\epsilon}$ yielding an overlapping Schwarz method, or we can add a multiplier that couples the two problems in the bulk domain $\Omega^{\epsilon}$, leading to the Arlequin method. A method of the former type was discussed on the discrete level in [16] and the Arlequin method was introduced in [17]. Define
$a_{k}(u^{k},v^{k}):=\int_{\Omega^{k}}\chi_{\epsilon,k}\nabla u^{k}\cdot\nabla v^{k}~{}\mbox{d}\Omega,$
where, for some $\chi$ with $10<\chi<$ 1,
(5.9) $\chi_{\epsilon,1}=\left\{\begin{array}[]{l}1\mbox{ in }\Omega\setminus\Omega^{\epsilon}\\ \chi\mbox{ in }\Omega^{\epsilon}\end{array}\right.\mbox{ and }\chi_{\epsilon,2}=\left\{\begin{array}[]{l}1\mbox{ in }\Omega\setminus\Omega^{\epsilon}\\ 1-\chi\mbox{ in }\Omega^{\epsilon}\end{array}\right.$
and (overlapping Schwarz method), with $\Gamma^{k}:=\partial\Omega^{\epsilon}\cap\Omega^{k}$, and $\mu:=\{\mu^{1},\mu^{2}\}$,
(5.10) $b_{k}(v,\mu):=(-1)^{k}\sum_{i=1}^{2}\left<\gamma_{\Gamma^{i}}v,\mu^{i}\right>_{\Gamma^{i}}$
or (Arlequin method)
(5.11) $b_{k}(v,\mu):=(-1)^{k}\int_{\Omega^{\epsilon}}\nabla v\cdot\nabla\mu+v\mu.$
For the overlapping Schwarz method we take $\mu^{k}\in\Lambda^{k}:=H^{-\frac{1}{2}}(\Gamma_{k})$, $k=1,2$ and define $\Lambda:=\Lambda_{1}\times\Lambda_{2}$. If we consider the Arlequin method on the other hand we define $\Lambda:=H^{1}(\Omega^{\epsilon})$. The weak formulation then takes the form, find $u^{k}\in V^{k}$, $i=1,2$ and $\lambda\in\Lambda$ such that
(5.12) $a_{1}(u^{1},v^{1})-b_{1}(v^{1},\lambda)$ $=(\chi_{\epsilon,1}f,v^{1})_{\Omega^{1}},\;\forall v^{1}\in V^{1},$
(5.13) $a_{2}(u^{2},v^{2})-b_{2}(v^{2},\lambda)$ $=(\chi_{\epsilon,2}f,v^{2})_{\Omega^{2}},\;\forall v^{2}\in V^{2},$
(5.14) $b_{1}(u^{1},\mu)+b_{2}(u^{2},\mu)$ $=0,\;\forall\mu\in\Lambda.$
The satisfaction of the infsup condition (5.10) and (5.11) is immediate by the definition of the forms in both cases. For (5.10) for the spaces $H^{\frac{1}{2}}(\Gamma^{k})$ and $H^{-\frac{1}{2}}(\Gamma^{k})$, $k=1,2$ and for (5.11) for the space $H^{1}(\Omega^{\epsilon})$ by the coercivity of the form defined in (5.11).
We readily verify that the local problem (5.12) is well-posed. If $\partial\Omega^{2}\cap\partial\Omega=\emptyset$, then the operator $A_{2}$ corresponding to the bilinear form $a_{2}$ has a non trivial kernel, consisting of constant functions on $\Omega^{2}$. It follows that the assumptions of our domain decomposition framework for stability are satisfied in both cases (5.10) and (5.11).
It remains to verify that the formulations are consistent with the original problem. First consider the Arlequin formulation using (5.11). Integrating by parts in (5.14) we see that there holds
(5.15) $\begin{array}[]{rcl}-\Delta(u^{1}-u^{2})+(u^{1}-u^{2})&=&0\mbox{ in }\Omega^{\epsilon},\\ \nu^{\epsilon}\cdot\nabla(u^{1}-u^{2})&=&0\mbox{ on }\partial\Omega^{\epsilon},\end{array}$
where $\nu^{\epsilon}$ denotes the outward pointing normal of the domain $\Omega^{\epsilon}$. This Neumann problem is well posed and hence $u^{1}=u^{2}$ in $\Omega^{\epsilon}$. The equations (5.12) and (5.13) then imply that
(5.16) $-\Delta u^{1}=f\mbox{ in }\Omega^{1}\setminus\Gamma^{2}\qquad\text{ and }\qquad-\Delta u^{2}=f\mbox{ in }\Omega^{2}\setminus\Gamma^{1},$
and
(5.17) $-\Delta\lambda+\lambda=0\mbox{ in }\Omega^{\epsilon}.$
On the boundary of $\partial\Omega^{\epsilon}$ there holds
$\nu^{\epsilon}\cdot\nabla u^{1}=\chi\nu^{\epsilon}\cdot\nabla u^{1}_{\epsilon}+\nu^{\epsilon}\cdot\nabla\lambda\mbox{ and }\nu^{\epsilon}\cdot\nabla\lambda=(1-\chi)\nu^{\epsilon}\cdot\nabla u^{2}\mbox{ on }\Gamma^{1},$
$\nu^{\epsilon}\cdot\nabla u^{2}=(1-\chi)\nu^{\epsilon}\cdot\nabla u^{2}_{\epsilon}+\nu^{\epsilon}\cdot\nabla\lambda\mbox{ and }\nu^{\epsilon}\cdot\nabla\lambda=\chi\nu^{\epsilon}\cdot\nabla u^{1}\mbox{ on }\Gamma^{2}.$
Here $u^{k}_{\epsilon}|_{\partial\Omega^{\epsilon}}:=\lim_{\varepsilon\rightarrow 0^{+}}u^{k}(x-\varepsilon n_{\epsilon})$ and $u^{k}|_{\partial\Omega^{\epsilon}}:=\lim_{\varepsilon\rightarrow 0^{+}}u^{k}(x+\varepsilon n_{\epsilon})$. Using that $u^{1}=u^{2}$ in $\Omega^{\epsilon}$ in the above interface conditions we immediately get
$\nu^{\epsilon}\cdot\nabla u^{1}=\nu^{\epsilon}\cdot\nabla u^{1}_{\epsilon}=\nu^{\epsilon}\cdot\nabla u^{2}_{\epsilon}\mbox{ on }\Gamma^{1}\mbox{ and }\nu^{\epsilon}\cdot\nabla u^{2}=\nu^{\epsilon}\cdot\nabla u^{2}_{\epsilon}=\nu^{\epsilon}\cdot\nabla u^{1}_{\epsilon}\mbox{ on }\Gamma^{2}$
as expected.
If instead (5.10) is used for the coupling in (5.12)-(5.14) the only difference to the above formulation is that we need a multiplier for each of the boundaries $\Gamma^{1}$ and $\Gamma^{2}$. $\lambda^{k}$ and $\mu^{k}$, $i=1,2$ are taken in $\Lambda^{k}:=H^{-\frac{1}{2}}(\Gamma^{k})$.
It follows as before that (5.16) holds. Instead of equation (5.15) we have that $-\Delta(u^{1}-u^{2})=0$ holds in $\Omega^{\epsilon}$ (this follows by multiplying (5.13) by $\chi/(1-\chi)$, subtracting (5.13) from (5.12) and testing with functions $v^{1}=v^{2}$ in $H^{1}_{0}(\Omega^{\epsilon})$). Since in this case, by the constraint equation (5.14), $u^{1}-u^{2}=0$ on $\partial\Omega^{\epsilon}$, we conclude also here that $u^{1}=u^{2}$ in $\Omega^{\epsilon}$. Finally on $\partial\Omega^{\epsilon}$ we now have
$\nu^{\epsilon}\cdot\nabla u^{1}=\chi\nu^{\epsilon}\cdot\nabla u^{1}_{\epsilon}+\lambda^{2}\quad\mbox{ and }\quad\lambda^{2}=(1-\chi)\nu^{\epsilon}\cdot\nabla u^{2}\quad\mbox{ on }\Gamma^{1},$
$\nu^{\epsilon}\cdot\nabla u^{2}=(1-\chi)\nu^{\epsilon}\cdot\nabla u^{2}_{\epsilon}+\lambda^{1}\quad\mbox{ and }\quad\lambda^{1}=\chi\nu^{\epsilon}\cdot\nabla u^{1}\quad\mbox{ on }\Gamma^{2}.$
Comparing the two cases we see that in both cases the solution of the Poisson problem is given by $u=u^{k}$ in $\Omega^{k}$. The multipliers of the two cases satisfy $\lambda^{k}=\nu^{\epsilon}\cdot\nabla\lambda$ on $\Gamma^{k}$, where $\lambda^{k}$ are the multipliers from the overlapping Schwarz method and $\lambda$ is the bulk multiplier from the Arlequin method. Hence the bulk multiplier of the Arlequin method can be reconstructed from the multiplier of the Schwarz method by solving (5.17) with $\lambda^{k}$ as Neumann data on $\Gamma^{k}$. These properties do not in general carry over to the discrete case, but will depend on how the different fields are discretized.
Remark 5.3.
When the form (5.11) is used the $\chi_{\epsilon,k}$ can be chosen as piecewise continuous functions summing up to $1$ in the interface zone. Then one can choose $a_{1}$ and $a_{2}$ to be different models and modify the weight to allow for a transition between the two models in the overlap zone.
Remark 5.4.
On the discrete level it is possible to choose $A_{h}^{+}$ as two completely different solvers for $a_{1}$ and $a_{2}$ using the above framework. For instance a computation using a particle method in $\Omega^{1}$ can in principle be coupled to a finite element solver in $\Omega^{2}$. A difficulty here is to make sense of the bulk version of the $b$ term (5.11) for methods not on variational form. Here it can be useful to discretize the multiplier using a structured mesh simplifying the interpolation problem and stabilize the multiplier as suggested in Section 3.1 to ensure that the infsup condition is satisfied.
5.3.2. Bulk-surface coupling
We consider a model for the interaction of concentrations in the bulk and on the surface, with a prescribed jump between the two. This is a special case of the model for proton transport in cell membranes introduced in [32]. For simplicity we let $\Omega$ be a smooth subset of $\mathbb{R}^{3}$ with boundary $\Gamma$ and outward pointing normal $\nu^{\Gamma}$. The model we consider takes the form
(5.18) $-\nabla\cdot(k^{B}\nabla u^{B})+u^{B}$ $=f^{B}\quad$ in $\text{in $\Omega$},$
(5.19) ${\sigma^{B}}u^{B}-{\sigma^{S}}u^{S}$ $=0\quad$ on $\text{on $\Gamma$},$
(5.20) $-\nabla_{\Gamma}\cdot(k^{S}\nabla_{\Gamma}u^{S})$ $=f^{S}-\nu^{\Gamma}\cdot k^{B}\nabla u^{B}\quad$ on $\text{on $\Gamma$}.$
Here $\nabla$ is the $\mathbb{R}^{3}$ gradient and $\nabla_{\Gamma}$ is the tangent gradient associated with $\Gamma$ defined by
(5.21) $\nabla_{\Gamma}=P_{\Gamma}\nabla$
with $P_{\Gamma}=P_{\Gamma}(x)$ the projection of $\mathbb{R}^{3}$ onto the tangent plane of $\Gamma$ at $x\in\Gamma$, defined by
(5.22) $P_{\Gamma}=I-n\otimes n.$
Further, ${\sigma_{B}}$, ${\sigma_{S}}$, $k_{B}$, and $k_{S}$ are positive constants, and $f_{B}:\Omega\rightarrow\mathbb{R}$ and $f_{S}:\Gamma\rightarrow\mathbb{R}$ are given square integrable functions. To define the weak form and the subproblems we define the spaces $V_{\Gamma}:=H^{1}(\Gamma)$, $V_{\Omega}:=H^{1}(\Omega)$ and $\Lambda:=H^{-\frac{1}{2}}(\Gamma)$, as well as the forms $a_{\Omega}:V_{\Omega}\times V_{\Omega}\to\mathbb{R}$ and $a_{\Gamma}:V_{\Gamma}\times V_{\Gamma}\to\mathbb{R}$
$a_{\Omega}(u^{B},v^{B}):=\int_{\Omega}\sigma^{B}(k^{B}\nabla u^{B}\cdot\nabla v^{B}+u^{B}v^{B})~{}\mbox{d}\Omega,\quad a_{\Gamma}(u^{S},v^{S}):=\int_{\Gamma}\sigma^{S}k^{S}\nabla_{\Gamma}u^{S}\cdot\nabla_{\Gamma}v^{S}~{}\mbox{d}\Gamma$
and the coupling form $b$,
$b_{1}(v,\mu):=\left<\mu,\sigma^{B}\gamma_{\Gamma}v\right>_{\Gamma},\,v\in H^{1}(\Omega),\mu\in\Lambda\mbox{ and }b_{2}(v,\mu):=-\left<\mu,\sigma^{S}v\right>_{\Gamma},\,v\in H^{\frac{1}{2}}(\Gamma),\mu\in\Lambda.$
The weak formulation may be written: find $u^{B},\,u^{S},\,\lambda\in V_{\Omega}\times V_{\Gamma}\times\Lambda$ such that
(5.23) $a_{\Omega}(u^{B},v^{B})-b_{1}(v^{B},\lambda)$ $=(\sigma^{B}f^{B},v^{B})_{\Omega}\mbox{ for all }\forall v^{B}\in V_{\Omega},$
(5.24) $a_{\Gamma}(u^{S},v^{S})-b_{2}(v^{S},\lambda)$ =$\sigma^{S}f^{S},v^{S}\right>_{\Gamma}\mbox{ for all }\forall v^{S}\in V_{\Gamma},$
(5.25) $b_{1}(u^{B},\mu)+b_{2}(u^{S},\mu)$ $=0\mbox{ for all }\forall\mu\in\Lambda.$
Here $\lambda:=\nu^{\Gamma}\cdot k^{B}\nabla u^{B}$ and the two subproblems are represented by the equations (5.23) and (5.24). Both the bulk equation (5.23) and the surface equation (5.24) are coercive, the surface equation (5.24) has a non-trivial kernel consisting of constant function s on $\Gamma$. The form $b$ is infsup stable, since it is defined by the duality pairing between $\Lambda$ and $V_{\Gamma}$, or the trace space of $V_{\Omega}$ on $\Gamma$. We conclude that the hypothesis of the abstract framework are satisfied.
Remark 5.5.
Also in this case the interpolation problem can be simplified by discretizing the multiplier space on a structured mesh when the spaces $V_{\Omega}$ and $V_{\Gamma}$ are discretized on unstructured meshes.
5.4. Branching fracture networks
In this final example we consider a toy problem for a fracture network in the form of three fractures whose boundaries meet in a common intersection $\Gamma$, a line in $\mathbb{R}^{d}$. Let $\Omega\equiv\cup_{k=1}^{3}\bar{\Omega}^{k}$ where the $\Omega^{k}$ are three different planar domains embedded in $\mathbb{R}^{3}$ with boundary $\partial\Omega^{k}$, outward pointing normal $\nu^{k}$ and common intersection $\Gamma$, i.e. $\overline{\Omega}^{k}\cap\overline{\Omega}^{j}=\Gamma$, $k,j=1,2,3$, $k\neq j$, where $\Gamma$. We now wish to solve the problem
(5.26) $-\nabla_{k}\cdot(a^{k}\nabla_{k}u^{k})$ $=f^{k}\mbox{ in }\Omega^{k},\,k=1,\ldots,3,$
(5.27) $u^{k}$ $=0\mbox{ on }\partial\Omega^{k}\setminus\Gamma,\,k=1,\ldots,3,$
[…]
Here the $a^{k}$ are positive constants, $\nabla_{k}$ denotes the differentiation operator in the plane $\Omega_{k}$ and $f^{k}\in L^{2}(\Omega^{k})$. We then introduce two Lagrange multiplier spaces $\Lambda_{1}(\Gamma)$ and $\Lambda_{2}(\Gamma)$, both taken as $H^{-\frac{1}{2}}_{00}(\Gamma)$, dual to $H^{\frac{1}{2}}_{00}(\Gamma)$ (if $M$ domains meet in $\Gamma$ we would need $M-1$ multipliers to impose the continuity (5.29) and the condition (5.28)). Also introduce the bulk spaces
$V_{k}:=H^{1}_{0}(\Omega^{k}):=\{v\in H^{1}(\Omega^{k}):v|_{\partial\Omega^{k}\setminus\Gamma}=0\}.$
Then we introduce the bilinear forms
$a_{k}(u^{k},v^{k}):=\int_{\Omega^{k}}a^{k}\nabla_{k}u^{k}\cdot\nabla_{k}v^{k}~{}\mbox{d}\Omega$
and
$b(v,\mu)=\left<\mu,\gamma_{\Gamma}v\right>_{\Gamma}.$
The weak formulation then takes the form: find $u^{k}\in V_{k}$, $k=1,2,3$ and $\lambda^{k}\in\Lambda_{k}$, $k=1,2$ such that
(5.30) $a_{1}(u^{1},v^{1})-b(\lambda^{1},v^{1})$ $=(f^{1},v^{1})_{\Omega^{1}}\quad\forall v^{1}\in V_{1}$
(5.31) $a_{2}(u^{2},v^{2})+b(\lambda^{1},v^{2})-b(\lambda^{2},v^{2})$ $=(f^{2},v^{2})_{\Omega^{1}}\quad\forall v^{2}\in V_{2},$
(5.32) $a_{3}(u^{3},v^{3})+b(\lambda^{2},v^{3})$ $=(f^{3},v^{3})_{\Omega^{1}}\quad\forall v^{3}\in V_{3},$
(5.33) $b(\mu^{1},u^{1}-u^{2})$ $=0\quad\forall\mu^{1}\in\Lambda_{1},$
(5.34) $b(\mu^{2},u^{2}-u^{3})$ $=0\quad\forall\mu^{2}\in\Lambda_{2}.$
The subproblems (5.30)–(5.32) are coercive with trivial kernel thanks to the Poincaré inequality. The infsup condition (2.4) of $b$ is ensured by the definition of the duality pairing. Remark that, in a more general setting where some fracture would couple to other fractures everywhere on its boundary, or has a Neumann condition, then the associated subproblem would have a kernel consisting of the piecewise constant functions.
To see that the condition (5.28) is satisfied we integrate by parts in the $a_{k}$ and use equation (5.26) to obtain,
$\nu^{1}\cdot(\mu^{1}\nabla_{1}u^{1})-\lambda^{1}=0,$
$\nu^{2}\cdot(\mu^{2}\nabla_{2}u^{2})-\lambda^{2}+\lambda^{1}=0,$
$\nu^{3}\cdot(\mu^{3}\nabla_{3}u^{3})+\lambda^{2}=0.$
Summing the three equations we obtain the condition (5.28). It follows that for this case also the assumptions of the abstract framework are satisfied.
Appendix A Construction and properties of the pseudo inverse
In this appendix we present some detail of the construction of a pseudo-inverse and we recall some of its properties, for more detail we refer to [3]. Throughout this section we let $X$, $Y$ denote two separable Hilbert space, and we let $H:X\to Y^{\prime}$ denote a bounded linear operator. We have the following propositions, which are not difficult to prove.
Proposition A.1.
$\mathfrak{R}(H)$ is finite dimensional iff $\ker(H)^{\perp}$ is, in which case the two have the same dimension.
Proposition A.2.
$H|_{\ker H^{\perp}}:\ker H^{\perp}\to\mathfrak{R}(H)$ is invertible. Moreover, letting $H^{-1}:\mathfrak{R}(H)\to\ker(H)^{\perp}$ denotes the inverse, we have the following identities
$H^{-1}Hx=x,\quad\forall x\in\ker(H)^{\perp},\qquad HH^{-1}y=y,\quad\forall y\in\mathfrak{R}(H).$
We can now construct a pseudo inverse as follows. We decompose $Y^{\prime}$ as $Y^{\prime}=\mathfrak{R}(H)\oplus\mathfrak{R}(H)^{\perp}$, and we define $H^{+}:Y^{\prime}\to X$ as
$\phi^{\prime}\in\mathfrak{R}(H),\ \phi^{\prime\prime}\in\mathfrak{R}(H)^{\perp}\quad\longrightarrow\quad H^{+}(\phi^{\prime}+\phi^{\prime\prime})=H^{-1}(\phi^{\prime}).$
We have the following proposition.
Proposition A.3.
$H^{+}$ is a pseudo inverse, that is
$HH^{+}Hx=Hx,\quad\forall x\in X,\qquad H^{+}HH^{+}\phi=H^{+}\phi,\quad\forall\phi\in Y^{\prime}.$
Moreover, we have that $\mathfrak{R}(H^{+})=\ker(H)^{\perp}$, and, letting $\Pi:X\to\ker(H)$ denote the orthogonal projection, we have
$H^{+}Hx+\Pi x=x,\qquad\forall x\in X.$
We remark that, in the situation where on either $X$ or $Y^{\prime}$, or both, different scalar products are available, while neither $\mathfrak{R}(H)$ nor $\ker(H)$ depend on which scalar product is chosen, the spaces $\ker(H)^{\perp}$ and $\mathfrak{R}(H)^{\perp}$, and consequently, the definition of $H^{+}$, do depend on such a choice. Choosing different scalar products results in different pseudo-inverses.
Appendix B Proof of Theorem 4.2
In order to prove Theorem 4.2 it is sufficient to prove that, for all $\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$ it holds that
$\frac{1}{C(V_{h},d,\sigma)}\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}\lesssim\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}.$
We start by observing that, letting ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$, we can split $B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}\in V_{h}$ as $B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}=z_{\widetilde{\varphi}}+w^{\perp}_{\widetilde{\varphi}}$, with $z_{\widetilde{\varphi}}\in Z$, and $w_{\widetilde{\varphi}}^{\perp}\in V_{h}$ $V$-orthogonal to $Z$. As $z\in\ker A=\ker A^{T}$, we have
$a(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}},B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}})=a(w^{\perp}_{\widetilde{\varphi}},w^{\perp}_{\widetilde{\varphi}}).$
The inf-sup condition (4.4) implies that $B_{\delta}$ is surjective. Then for some $w^{\prime}\in V_{h}$, it holds that $\mathcal{K}{\widetilde{\varphi}}=B_{\delta}w^{\prime}$. Then we can write
$\mathcal{K}{\widetilde{\varphi}}=B_{\delta}w^{\prime}=B_{\delta}B_{\delta}^{+}B_{\delta}w^{\prime}=B_{\delta}B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}.$
Let now $\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$. We have
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\gtrsim\|\widehat{\lambda}\|_{\Lambda}\gtrsim\sup_{w\in V_{h}}\frac{b(w,\widehat{\lambda})}{\|w\|_{V}}{\geq}\sup_{w\in V_{h}\cap Z^{\perp}}\frac{b(w,\widehat{\lambda})}{\|w\|_{V}}\gtrsim\sup_{w\in V_{h}\cap Z^{\perp}}\frac{b(w,\widehat{\lambda})}{\sqrt{a(w,w)}}$
$\geq\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{b(w^{\perp}_{\widetilde{\varphi}},\widehat{\lambda})}{\sqrt{a(w^{\perp}_{\widetilde{\varphi}},w^{\perp}_{\widetilde{\varphi}})}}=\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{b(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}}-z_{\widetilde{\varphi}},\widehat{\lambda})}{\sqrt{a(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}},B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}})}}=\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{b(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}},\widehat{\lambda})}{\sqrt{a(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}},B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}})}}$
$=\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{a(B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}},B_{\delta}^{+}\mathcal{K}{\widetilde{\varphi}})}}=\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}$
and we have the upper bound.
Let us now prove the lower bound. Recalling that $B_{\delta}^{+}B_{\delta}=\Pi_{d}$ (see (4.5)), we have
(B.1) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\widehat{\lambda}\|_{\Lambda}\lesssim\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|v\|_{V}}=\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|\Pi_{d}v\|_{V}}\frac{\|\Pi_{d}v\|_{V}}{\|v\|_{V}}\\ \leq\|\Pi_{d}\|_{V\to V}\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|\Pi_{d}v\|_{V}}=\|\Pi_{d}\|_{V\to V}\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}v\|_{V}}.$
We also have
$b(w,\widehat{\lambda})=\langle B_{\delta}w,\widehat{\lambda}\rangle=\langle B_{\delta}B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}\rangle=b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}),$
whence
(B.2) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\Pi_{d}\|_{V\to V}\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
Now, for $z\in Z$ arbitrary, we can write
(B.3) $\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}=\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}(w+z),\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
The idea is now to take $z$ such that $B_{\delta}^{+}B_{\delta}(w+z)\in\mathfrak{R}(\mathcal{K})$. To show that this is possible we observe that, by the definition of $\mathcal{K}$, given $\phi\in\Lambda_{\delta}^{\prime}$ we have that $\zeta=\phi-\mathcal{K}([\phi])\in\ker\mathcal{R}$. Then, for all $\lambda\in\Lambda_{\delta}$
$0=\langle\zeta,\Pi_{\sigma}\lambda\rangle=\langle\Pi_{\sigma}^{T}\zeta,\lambda\rangle,$
that is, $\zeta\in\ker\Pi_{\sigma}^{T}$. We have the following lemma.
Lemma B.1.
Letting $G_{\delta}:Z\to\Lambda_{\delta}^{\prime}$ denote the restriction of $B_{\delta}$ to $Z$, it holds that $\ker\Pi^{T}_{\sigma}=\mathfrak{R}(G_{\delta})$.
Proof.
If $\phi=G_{\delta}z=B_{\delta}z$ then $\Pi_{\sigma}^{T}\phi$ satisfies
$\langle\Pi_{\sigma}^{T}\phi,\lambda\rangle=\langle\phi,\Pi_{\sigma}\lambda\rangle=\langle B_{\delta}z,\Pi_{\sigma}\lambda\rangle=0$
and we then have $\mathfrak{R}(G_{\delta})\subseteq\ker\Pi_{\sigma}^{T}$. Let now $\phi\in\ker\Pi_{\sigma}^{T}$. On the one hand we can split $\phi$ as $\phi=G_{\delta}z+\phi^{\perp}$ with $\sigma^{*}(\phi^{\perp},G_{\delta}z^{\prime})=0$ for all $z^{\prime}\in Z$. On the other hand we have, for all $\lambda\in\Lambda_{\delta}$,
$0=\langle\Pi_{\sigma}^{T}\phi,\lambda\rangle=\langle\phi,\Pi_{\sigma}\lambda\rangle=\langle G_{\delta}z+\phi^{\perp},\Pi_{\sigma}\lambda\rangle=\langle\phi^{\perp},\Pi_{\sigma}\lambda\rangle.$
Then we have that ϕsuperscriptitalic-$\phi^{\perp}$ simultaneously verifies
$\sigma^{*}(\phi^{\perp},G_{\delta}z)=0,\ \forall z\in Z,\qquad\text{ and }\qquad\langle\phi^{\perp},\widehat{\lambda}\rangle=0,\ \forall\widehat{\lambda}\in\widehat{\Lambda}_{\delta}.$
We claim that this implies that $\phi^{\perp}=0$, which, in turn, implies that $\phi\in\mathfrak{R}(G_{\delta})$. To prove the claim we observe that
$\sigma^{*}(\phi^{\perp},G_{\delta}z)=\langle\Sigma^{-1}\phi^{\perp},G_{\delta}z\rangle=0,$
which implies $\mu=\Sigma^{-1}\phi^{\perp}\in\widehat{\Lambda}_{\delta}$. But we have that
$\sigma(\mu,\widehat{\lambda})=\langle\Sigma\Sigma^{-1}\phi^{\perp},\widehat{\lambda}\rangle=0,\ \forall\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$
and this implies $\mu=0$ and, consequently $\phi^{\perp}=0$, which, in turn, implies $\phi\in\mathfrak{R}(G_{\delta})$. ∎
We conclude that $\ker\mathcal{R}=\mathfrak{R}(G_{\delta})$, so that for all $\phi$ we have, for some $z\in Z$, $\phi-\mathcal{K}[\phi]=G_{\delta}z$. Then we can choose $z^{*}\in Z$ as
$z^{*}=\mathop{\text{argmin}}_{z:\ G_{\delta}z=(\mathcal{K}-1)B_{\delta}w}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
We have, with ${\widetilde{\varphi}}=[B_{\delta}(w+z^{*})]$,
(B.4) $\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle=b(B_{\delta}^{+}B_{\delta}(w+z^{*}),\widehat{\lambda})=b(B_{\delta}^{+}B_{\delta}w+z^{*},\widehat{\lambda})=b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}),$
as well as
(B.5) $\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})=m(\mathcal{K}[B_{\delta}(w+z^{*})],\mathcal{K}[B_{\delta}(w+z^{*}))=m(\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}),\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}))\\ =a(B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}),B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}))\lesssim\|B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*})\|_{V}^{2}=\|B_{\delta}^{+}B_{\delta}(w+z^{*})\|^{2}_{V},$
where we could write $\Pi_{\sigma}^{T}B_{\delta}(w+z^{*})=B_{\delta}(w+z^{*})$ thanks to our choice of $z^{*}$. Plugging (B.4) and (B.5) into (B.3), and the resulting bound into (B.2) we obtain
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\Pi_{d}\|_{V\to V}\left(\sup_{w\in V_{h}}\min_{\begin{subarray}{c}z\in Z\\ G_{\delta}z=(\Pi_{\sigma}^{T}-1)B_{\delta}w\end{subarray}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}\right)\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}.$
It remains to bound
$\sup_{w\in V_{h}}\min_{\begin{subarray}{c}z\in Z\\ G_{\delta}z=(\Pi_{\sigma}^{T}-1)B_{\delta}w\end{subarray}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
We have
$\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}\leq\|B_{\delta}^{+}B_{\delta}w\|_{V}+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}w\|_{V}\\ =\|B_{\delta}^{+}B_{\delta}w\|_{V}+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}B_{\delta}^{+}B_{\delta}w\|\\ \lesssim\left(1+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}\|_{V\to V}\right)\|B_{\delta}^{+}B_{\delta}w\|_{V}.$
This finally yields
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim C(V_{h},d,\sigma)\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}.$
Acknowledment
EB was partially funded by the EPSRC project EP/W007460/1, Integrated Simulation at the Exascale: coupling, synthesis and performance. SB is member of the Gruppo Nazionale di Calcolo Scientifico – Istituto Nazionale di Alta Matematica (GNCS–INdAM).